A Complete Bibliography of Publications in Canadian Journal of Mathematics = Journal canadien de

mathématiques for the decade 1960–1969

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.08

Title word cross-reference

(0, 1) [84, 164, 227, 207, 267, 418]. (1, 1) [74]. (C, 1) [442]. (C, γ, μ) [369]. (f, d_{n}) [388, 993]. (x, y, z) = (y, z, x) [778]. (Z, p) [330]. 0 [863, 934]. 1 [789]. 2 [978, 459, 422, 878, 802]. 3 [245]. 4 × n [660]. 5 [626]. 52 [987]. [5] [452]. [F, d_{n}] [993, 60, 240]. [J, f(x)] [366]. ABA [138]. α [161, 606]. C(X) [725]. C(H) [154]. D [925]. P [591]. D [520, 600]. D(X) [906]. e [712]. E^{n} [819]. F_{5} [864]. C [921]. D_{n} [855]. F [909]. H [420]. H^{1}(R) [590]. H^{∞} [412]. H^{p} [861, 327]. I [602, 741]. II_{1} [25]. J [708, 685]. K [429, 283, 404, 409, 824, 36]. L(r, t) [560]. L^{p} [916, 914]. L^{p}(−∞, ∞) [106]. Ł_{a}^{p} [108]. L_{p} [562]. m [753, 359]. M(phi) [646]. M_{3}^{8} [467]. modulo1 [632]. μ(z, β, α)

[955]. N [515, 979, 385, 177, 647, 198, 866, 624, 473, 961, 663]. n + 1 [663]. nheta [789]. P [520, 619, 797, 12, 612, 764]. p = 4q + 1 [619]. P^{−1}XP = dg(λ_{1}, ⋯, λ_{n}) = D [520]. p^{3} [248]. p_{−k}(n) [699, 868, 898]. P AQ − QAP [825]. PG(3, q) [587]. PG(N, q^{2}) [553]. π [471, 684]. P Q [748]. PSL(2, p) [834, 875]. PSp_{4}, (3) [650]. q [619]. r [671]. R^{4} [410]. R_{n} [784]. S [770]. S_{2} [736]. σ [929]. σ_{k}(n), k > 1 [947]. Sn [278]. sum_{i=1}^{∞}A_{i}e^{B_{i}}X = X [331]. T_{2} [373]. v [633, 614]. V_{m} [196]. V_{n} [196]. W′′ + pW = 0 [129]. X [520]. X, Y [922]. z [570]. z^{−1} [570].

-abelian [979]. -adic [12]. -algebras [929, 748]. -anneau [708]. -ANR [624]. -ball [198]. -bisimple [741, 602]. -classes [925]. -colored [36]. -coloured [824]. -complements [878]. -convex [753]. -cube [385]. -cyclic [764]. -dimensional [671, 459, 863, 934]. -dimensionalen [671]. -extremally [283]. -functions [770]. -generated [161]. -graphs [961]. -groups [612, 600, 373, 736]. -homogeneity [369]. -like [591]. -matrices [84, 164, 227, 207, 267, 418]. -method [388]. -multiplication [633]. -normal [404]. -product [429]. -quota [359]. -regular [471]. -ring [708]. -rings [614, 685]. -saturated [409]. -Schauder [712]. -space [177, 663]. -spaces [914]. -sphere [245]. -spiral [606]. -stable [797]. -subgroups [978, 422]. -surfaces [420].

2-betweenness [533]. 2-dimensional [729, 771]. 2-metric [533].

A. [441]. abelian [979, 157, 259, 422, 335, 814, 134, 683, 901, 886]. abélienne [20, 194]. Absolute [861, 361, 905, 707, 938]. Abstract [977, 72, 187]. accessory [209]. acting [962]. acyclic [596]. Adams [714]. addendum [742, 469, 740]. addition [1002]. Additive [562]. adic [12]. Adjoint [901, 59, 755, 999]. adjoints [13]. admissible [143]. admit [9]. admitting [818]. Affine [430, 306, 406, 850, 264, 146, 325, 962]. affines [962]. agissant [962]. albelian [25]. algebra [738, 647, 251, 397, 576, 577, 1008, 276, 928, 638, 144, 675, 830, 158, 467, 573, 127]. algebraic [840, 774]. Algebras [984, 30, 881, 120, 121, 582, 809, 492, 169, 835, 721, 753, 848, 483, 395, 455, 972, 625, 847, 899, 419, 800, 747, 748, 620, 466, 757, 1014, 718, 74, 44, 915, 594, 143, 113, 1001, 727, 145, 273, 345, 767, 967, 929, 717, 296, 393, 467, 774, 803, 216, 81, 178, 913, 970, 895, 943]. Algorithm [703, 426, 187]. algorithmic [311, 387]. Almost [363, 281, 806, 107, 892, 691, 358]. Almost-bounded [281]. almost-Hermitian [358]. alternating [319]. Alternative [512, 396, 350]. amalgamated [446]. ambiguous [281]. among [190]. analogue [312, 849]. analogues [722, 441]. analysis [770, 15]. Analytic [148, 239, 280, 118, 355, 550, 687, 974, 779, 971, 67, 393, 7]. analyticity [117, 118]. André [879]. Angular [149, 235]. Anneau [708, 10]. Anneaux [356]. annihilator [492]. annuli [355]. ANR [624]. anti [296, 393]. anti-commutative [296, 393]. Anticommuting [114]. Antisymmetrical [669]. any [219]. applications [71, 68, 656, 1008, 15, 243]. approach [556, 150]. Approximate [160]. Approximation [728, 922, 555, 57]. approximations [574]. arbitrary [470]. arc [19]. Archimedean [954]. Arcs [291, 805, 344, 465, 598, 663, 751, 819]. area [615, 346]. arising [251]. Arithmetic [72, 844, 161, 449, 571, 882, 429, 950]. Arithmetical [35]. arithmétique [571, 950]. arrays [329]. Artinian [652]. Associated [795, 452, 976, 905, 904, 255, 517]. associative [396, 647, 605, 74, 345, 767, 216, 894, 895]. asymmetrical [705]. Asymptotic [176, 105, 377, 692, 616, 891, 224, 955]. asymptotically [376]. automata [347, 713]. Automorphism [970, 9, 818, 559, 403, 309, 793]. automorphismes [848]. Automorphisms [55, 239, 982, 942, 219]. average [443]. averages [232]. Axes [665]. axiom [853]. axiomatic [781]. Axioms [707, 938].

b. [780, 956]. balanced [16, 321, 348, 253, 1012]. Balázs [849]. ball [198]. Banach [974, 106, 915, 976, 273, 799, 909, 803, 901, 105, 178]. base [583]. based [133]. bases [229, 357, 976, 258, 564, 645, 712]. basis [534]. be [304]. behaviour [891, 176, 904, 946, 224, 955]. Beltrami [399]. Bergman [203, 563]. Bernoulli [170]. Bernstein [880, 282, 448]. Bers [399]. Betti [518]. between [859, 860, 21, 672, 399, 544, 290, 817, 5, 322, 470, 918, 551]. betweenness [533]. Beurling [374]. bi [82]. bi-coloured [82]. bicolourable [213]. Bicontinuous [551]. biharmonic [437]. bilateral [4]. bilinear [286, 491]. binary [228]. Biorthogonal [914, 976]. bipartite [964]. bipartitional [251]. Bipositive [419]. Birkhoff [592]. Bisecants [476]. Bisimple [694, 602, 741]. BK [911]. BK-spaces [911]. Blaschke [149, 903, 543, 910]. Block [73, 16, 462, 321, 348, 253, 1012]. bodies [386, 101, 657]. Bol [998]. Boolean [158, 232]. Borel [926]. Borel-type [926]. Borsuk [486]. bound [920]. Boundary [474, 548, 115, 103, 550, 756, 358, 465, 563, 1004, 523, 827]. boundary-value [523, 827]. bounded [981, 19, 281, 230, 583, 728, 288, 308, 1004, 909, 537]. boundedness [140]. Bounds [563, 272, 893, 629]. bracket [845]. branch [918]. Brauer [920]. Brualdi [792]. Brunn [101]. bundle [499]. bundles [610, 499].

cacti [992]. calculus [752]. can [601]. cancellation [365]. Canonical [39]. Caradus [909]. cardinalities [268]. case [687, 545, 67]. Casimir [1007]. categorical [841]. Catégories [750, 862, 134, 750, 413]. category [1008, 924]. Cauchy [360, 688, 664]. Cayley [4, 266, 216]. Cech [936]. Cell [648, 294, 124]. census [347, 155, 125, 184, 214]. centered [535]. central [195, 848, 847, 378]. centralizer [895]. centralizers [422, 878]. centrally [677]. centres [682, 51]. Cercles [891]. Certain [4, 80, 410, 738, 865, 222, 444, 568, 742, 154, 108, 183, 256, 725, 351, 333, 469, 452, 371, 228, 438, 439, 440, 624, 3, 443, 293, 159, 310, 969, 870, 665, 188, 39, 300, 128, 680]. certaines [108]. certains [410]. Cesàro [762, 905]. chain [333, 469]. chainable [107, 508, 883]. chains [24, 76]. character [246]. characteristic [42, 376, 797, 823, 145, 802]. Characterization [228, 855, 533, 883, 314, 477, 739, 650, 371, 373, 928, 424]. Characterizations [850, 417]. characterize [790]. Characters [204, 838, 820, 836]. Chevalley [942]. chords [344]. chromatic [48, 871, 534]. chromatiques [871]. chronogeometry [670]. circle [132, 584, 370, 168]. circle-like [132, 584]. circles [672]. circuits [372, 418]. circular [463]. circularly [508]. class [30, 80, 384, 588, 763, 646, 462, 911, 405, 756, 207, 59, 91, 308, 755, 934, 174, 22, 267, 350, 909, 912, 833, 341, 494]. Classes [455, 495, 227, 66, 539, 611, 732, 925]. classical [116]. Classification [575, 979, 995, 676, 727]. Clifford [441, 127]. climbing [524]. Closed [743, 19, 760, 807, 551, 179]. closure [255]. closures [954]. Cluster [198]. co [532, 342]. co-ordinatizing [342]. co-rational [532]. coarseness [964, 775]. cobounding [46]. Cocommutative [594]. codes [206]. codimension [918]. Coefficients [1004, 90]. coercive [756]. cohomologie [20, 194]. cohomology [20, 194, 481]. cohomotopy [923]. coincide [116]. collections [476]. Collineation [653, 879, 306, 864]. Collineations [264, 323, 662, 831, 504]. colored [36]. colour [667]. coloured [220, 824, 82, 122, 271]. Colouring [402, 744]. column [711]. combinatorial [444, 312, 438, 439, 440, 792, 770, 15, 744]. comma [206]. comma-free [206]. Common [622]. Commutative [767, 380, 53, 365, 44, 915, 192, 296, 393]. Commutativity [693, 843, 685, 22]. Commutators [407, 552, 969, 709, 511, 290, 23, 389, 94]. commuting [900, 622]. compact [135, 1011, 631, 962, 397, 490, 840, 828, 873, 934, 551]. compacta [591]. compactifications [497]. compactly [38]. compactness [279]. compacts [796]. comparability [314]. comparison [115, 649]. compatible [463]. complementary [845]. complementation [447]. Complemented [273, 765, 178, 769]. complements [878]. Complete [285, 815, 853, 421, 964, 952, 890, 775, 702, 603, 573, 480]. completely [890]. completeness [1009]. completion [404]. Completions [265, 352, 338, 862]. complex [697, 880, 75, 240, 223, 187, 909]. complex-valued [223]. Complexes [573, 294, 449]. Composition [730, 16]. compositions [509, 813]. Compounds [307]. concept [609]. Concerning [584, 301, 819, 981, 876, 757, 1014, 772, 993, 971, 292, 366, 664]. condition [121, 810, 785, 539]. conditions [103, 684, 333, 469, 304, 209, 870]. cones [459, 976]. configurations [254, 438, 439, 440, 348]. congruence [39, 123]. congruences [699, 868, 898, 217, 597, 493, 821, 293]. Conical [277]. conics [662]. Conilpotency [924]. conjecture [17]. conjugate [238, 147, 905, 904, 732]. Conjugates [580]. connected [463, 931, 1011, 228, 764]. Connections [315, 899, 5]. Connectivity [167, 566, 50, 49, 689]. consecutive [289, 205, 628]. consequence [1014]. consequences [757]. constant [671, 427]. constants [88, 185]. Construction [68, 939, 526, 968, 310, 734, 17, 183, 963, 403, 780, 956, 145, 564, 456]. contained [372, 328]. containing [653]. containment [139]. contains [19]. continua [132, 107, 508, 883, 584, 819]. continuation [380, 67]. continued [791, 27]. continuity [361, 119, 160, 480]. Continuous [489, 61, 119, 28, 151, 568, 742, 154, 126, 400, 223, 644, 819, 54, 902]. continuum [19]. contractible [729, 771]. contraction [547]. contractive [829]. contribution [670]. Contributions [124]. Convergence [791, 500, 861, 482, 810, 66, 503, 806]. converse [182, 547]. Convex [706, 458, 613, 753, 386, 962, 344, 101, 657, 261, 989, 51, 326, 950, 78, 525, 580, 664, 565, 678, 202, 334]. convexe [962]. convexes [950]. Convexity [414, 717]. convolution [419, 394]. Convolutions [286]. coordinate [645]. coprime [820]. corpoidal [10, 10]. Correction [844, 153, 387, 903]. Corrections [303]. correlations [662]. Correspondences [838, 542]. corresponding [219]. Corrigendum [898, 742, 983, 1014, 956, 740, 1012, 1013]. cosine [6]. Countable [497, 927]. countable-dimensional [927]. countably [948]. Counting [122, 271, 375]. couples [684, 936]. cover [187]. Covering [611, 715, 321, 262]. coverings [177]. covers [627]. Criteria [293, 579, 983]. criterion [867, 975]. crystallographic [676]. Csima [703]. cube [385]. current [280]. curvature [671]. curve [19, 613, 344, 334]. curves [489, 51, 663, 196, 902]. Cycles [689]. cyclic [96, 403, 764, 751, 40, 109].

D.C.C. [116]. dans [871].

Darstellungstheorie [754]. decision [244]. Decomposition [863, 76, 237, 930, 308, 782, 589, 593, 383]. decompositions [354, 952, 531, 723, 819]. decreasing [257, 673, 79]. Dedekind [11, 258]. defining [875]. definite [986]. degenerate [360, 863, 819]. degree [619, 288, 767, 957, 958]. degrees [65, 856]. deleted [729, 771]. Demushkin [575]. derivate [839]. derivates [705]. derivation [81]. derivations [881]. Derivative [339, 866]. Derivatives [583, 68, 149, 673]. derived [851, 87]. Desarguesian [306]. descendent [501]. description [412]. design [462, 73]. designs [16, 321, 780, 956, 348, 253, 1012]. determinant [709]. Determination [444, 749]. Determining [268]. developable [417]. deviation [503]. diagonal [659]. diameter [764]. Diameters [320]. difference [190, 495, 175, 335, 814, 581, 391, 199, 298, 349]. Differences [673, 195]. different [713]. Differentiable [1006, 615, 705, 644, 663]. differential [618, 28, 151, 399, 687, 405, 756, 760, 104, 579, 983, 59, 91, 893, 866, 176, 563, 649, 629, 150, 414]. differentiation [539, 242, 277, 130, 160]. digits [322, 33]. digraph [764]. digraphs [522, 669]. dimension [696, 110, 727]. dimensional [671, 483, 459, 98, 854, 863, 934, 729, 771, 294, 325, 346, 927]. dimensionalen [671]. dimensions [759, 515, 473]. diminishing [269]. diophantine [34, 57, 398]. Direct [639, 774, 454, 890, 249, 446]. Directed [627, 460, 507]. Dirichlet [876, 2]. disconjugacy [866]. disconnected [283]. Discrete [37, 865, 997, 918]. discriminant [64]. disjoint [826, 328, 587, 418, 197]. disk [367]. disks [521]. dissections [367]. distance [672]. Distinct [1005, 502]. distortion [355]. distributed [415, 806]. distribution [362, 616, 186, 935, 789, 5, 632]. distributions [567]. Distributive [86, 466, 765, 856]. divide [344]. divided [195]. divisibility [528, 908]. divisible [201]. division [193, 994]. divisor [274]. Domain [437, 906, 812, 53]. domains [785, 11, 201, 110, 563, 540]. dominant [727]. Double [569, 384, 895]. double-centralizer [895]. doublement [336]. doubly [137, 513]. Dual [250, 181, 101, 490, 840, 609, 551]. duality [312, 990]. duals [720]. due [679]. d’un [708, 10]. dynamical [65].

edge [200, 961]. edges [353, 197]. eight [546]. Elementary [471, 678]. elements [651, 852, 647, 126, 863, 240, 45, 819]. ellipses [816]. elliptic [548, 399, 405, 756, 737, 649, 523, 827]. elliptic-parabolic [649]. embeddability [952]. Embedding [132, 259, 591, 191, 544, 488]. Embeddings [508]. Endomorphism [776]. endomorphisms [380, 682]. Engel [493, 821]. ensemble [962]. ensembles [871]. Entire [376, 697, 509, 813, 549, 869, 1013, 243]. Entropies [230]. Enumeration [241, 367, 213, 522, 695, 48, 317, 794, 82, 578]. equal [344, 294, 917, 322]. equality [545]. equally [165]. equation [318, 12, 472, 6, 688, 150]. equations [399, 687, 181, 405, 760, 104, 579, 983, 338, 772, 866, 250, 563, 649, 491, 92, 313, 105, 171, 414]. equicontinuous [357]. equilateral [599]. Equilibrium [596]. equivalence [31, 223]. equivalences [696]. Equivalent [486, 304, 411]. Errata [741]. error [574]. espace [108]. espaces [796]. essential [893, 557]. Estimates [756, 574, 366]. Euclidean [432, 676, 1001, 679, 179]. Euler [17, 507, 162, 337, 832]. even [28, 944]. events [713]. exact [684]. exceptional [654]. existence [626, 253, 1012, 446]. Expansion [644, 103]. expansions [515]. exponent [493, 821, 1010]. exponential [398]. exponents [931]. expression [63, 300]. Extension [284, 21, 532, 810, 637, 609, 350, 751, 529]. Extensions [682, 536, 602, 741, 965, 835, 608, 435, 1010, 466, 258, 504]. exterior [276, 370, 675, 168, 130]. Extrema [506]. Extremal [370, 168, 961, 613]. extremally [283]. Extreme [590, 779, 799, 456]. extremization [711]. extremum [31, 173].

faces [261, 458, 677, 346]. factor [948]. factorization [737, 999, 471]. factors [552, 861, 985, 25]. falsity [17]. families [260, 489, 902]. famille [336]. Federer [936]. Fermat [236]. few [9, 836]. fiberings [1006, 859, 860, 210]. Fibonacci [473]. fibre [610]. field [965, 87, 491, 912, 171, 1005]. fields [500, 728, 998, 808, 403, 358, 258, 968, 133, 529]. filter [412]. final [774]. Finite [948, 890, 759, 9, 422, 878, 811, 163, 305, 732, 666, 957, 958, 1006, 8, 246, 889, 553, 157, 876, 483, 103, 284, 476, 306, 138, 347, 464, 406, 758, 650, 850, 783, 373, 598, 842, 50, 76, 491, 133, 348, 430, 383, 864, 959, 55, 56, 313, 1005]. finite-dimensional [483]. finitely [441, 174, 969]. Finiteness [638]. Finsler [790, 131]. first [922]. Fixed [820, 71, 592, 749, 502, 900, 818, 842, 622]. fixed-point [900]. fixed-point-free [818]. flag [306]. flag-transitive [306]. flat [530, 130]. Flowers [382]. fonctions [950]. form [434, 825, 63, 498, 138, 296, 123]. Formal [413, 331]. Formally [435]. forms [169, 229, 977, 456, 130, 39, 529, 208, 219]. formula [112, 442, 947]. formulas [35, 302, 142]. formulations [486]. Foundations [65]. Four [696, 344, 405, 816, 291, 767, 505, 667, 346]. four-colour [667]. four-dimensional [346]. four-person [505]. Fourier [63, 442, 66, 905, 904, 946, 644]. Fourth [115, 893]. fourth-order [893]. Fractional [181, 996]. fractions [708, 791, 27, 537]. framal [466]. Frattini [889]. Fredholm [738, 852]. Free [831, 698, 272, 748, 86, 747, 818, 206, 85, 163, 201, 297]. freedom [65]. freely [161]. French [571, 410, 871, 20, 194, 108, 29, 708, 962, 950, 750, 10, 356, 2]. Frobenius [913]. Fubini [351]. function [697, 30, 906, 398, 980, 404, 251, 154, 152, 583, 845, 274, 808, 933, 136, 839, 717, 869, 1013, 943]. functional [6]. functionals [495, 562, 803]. functions [239, 339, 257, 361, 119, 981, 475, 588, 763, 618, 118, 882, 235, 281, 855, 876, 238, 198, 230, 377, 692, 233, 455, 376, 212, 974, 215, 779, 705, 429, 900, 509, 737, 813, 515, 620, 400, 884, 89, 873, 950, 308, 465, 606, 971, 807, 897, 3, 61, 203, 928, 526, 370, 951, 223, 5, 549, 770, 520, 1004, 580, 799, 168, 644, 833, 984, 247, 33, 986, 704, 243]. fundamental [78]. Further [17, 461].

G. [679]. Galois [899, 133, 870]. game [505]. games [462, 73, 359]. Gap [478]. General [116, 484, 588, 763, 340, 607, 756, 269, 505, 717, 171]. general-sum [505]. Generalization [919, 473, 857, 581, 295, 912, 131, 671]. Generalizations [547]. generalize [467]. Generalized [889, 997, 701, 1007, 657, 45, 329, 738, 475, 68, 190, 156, 303, 583, 879, 640, 515, 338, 623, 975, 292, 762, 688, 49, 951, 858, 846, 446, 349, 315]. generated [617, 98, 272, 161, 441, 996, 38, 969, 468]. Generating [704]. Generation [802, 378]. Generators [141]. genetic [680]. genus [385, 518]. geodesics [256]. Geometric [680]. geometrical [675, 150, 752]. Geometrie [651, 965]. geometries [342]. geometry [651, 965, 656, 133, 707, 938, 325, 679, 781, 131]. German [651, 965, 671]. Gibbs [60, 185]. girth [626, 944]. girths [546]. given [18, 1008, 601, 203, 309, 695, 732, 793, 716]. global [912]. Goldie [941]. Gotlieb [703]. Gram [678]. Graph [93, 744, 463, 421, 964, 518, 372, 451, 656, 534, 775, 460, 406, 488, 702, 189, 418, 197]. Graph-colouring [744]. graphes [871]. Graphs [944, 425, 48, 510, 952, 931, 546, 626, 871, 220, 183, 353, 314, 409, 213, 642, 558, 1008, 167, 371, 824, 320, 82, 841, 390, 50, 76, 237, 507, 695, 411, 36, 162, 383, 337, 961, 793, 541, 603, 270, 689, 122, 271]. Grassmann [888, 208]. greatest [716]. Green’s [618]. Group [576, 577, 651, 157, 937, 395, 252, 397, 215, 797, 460, 490, 840, 650, 335, 814, 945, 403, 441, 787, 38, 311, 387, 309, 392, 967, 294, 870, 886, 325, 887, 973, 793, 127, 219, 123]. groupes [571, 29, 336]. Groups [978, 452, 288, 836, 996, 917, 979, 96, 619, 571, 80, 8, 652, 246, 834, 889, 384, 960, 948, 42, 157, 954, 259, 380, 907, 248, 75, 98, 653, 29, 285, 631, 397, 139, 306, 879, 759, 640, 493, 838, 140, 9, 138, 422, 878, 37, 818, 821, 724, 930, 612, 811, 758, 942, 739, 559, 820, 95, 249, 963, 305, 575, 854, 62, 373, 736, 676, 174, 908, 683, 732, 802, 319, 846, 310, 969, 966, 478, 864, 886, 55, 56, 141, 40, 109, 957, 958, 600, 970, 875]. growth [377, 692, 648, 910, 5, 124]. Gruppenelemente [651].

H [441]. Hadamard [987, 190, 329]. Half [559, 756]. half-space [756]. Half-transitive [559]. Hall [251]. halo [539]. Hamiltonian [155]. Hardy [623, 295]. harmonic [630, 203]. Hausdorff [88, 366, 102]. having [377, 692, 435, 913]. heights [84]. Heine [561]. Hereditary [396, 881, 734]. Hermite [224]. Hermitian [553, 892, 358]. Herstein [991]. Hessenberg [679]. hexagons [256]. hide [601]. higher [651, 682, 290, 324, 170, 389]. Hilbert [621, 407, 544, 6, 414]. Hjelmslev [851, 284]. Höhensatz [651]. Hölder [61, 857]. Holomorphic [717, 981, 281, 984]. homeomorphism [829]. homeomorphisms [21, 691]. homeotopy [294]. homogeneity [369]. Homogeneous [107, 19, 313, 499]. homological [379]. Homology [729, 771, 886]. homomorphic [716]. homomorphism [276]. homomorphisms [724, 412]. homotheties [828]. homotopic [630]. Homotopical [896, 481]. Homotopy [316, 77, 134, 966, 595, 843, 766, 817]. homotopy-commutativity [843]. Hopf [148, 594]. Hughes [424]. Hyperbolic [394, 674, 67]. hyperdegrees [856]. hyperfinite [985]. hypergeometric [4]. hyperhomology [739]. hypernilpotent [605]. hyperspaces [805]. hypersurface [790]. hypersurfaces [604].

Ideal [354, 906, 53, 218, 531, 144]. Ideals [263, 554, 932, 53, 365, 166, 343, 773, 915, 874, 540, 959, 551]. Idempotent [874]. Identities [345, 4, 120, 733, 915, 767]. II [860, 809, 889, 763, 607, 634, 635, 362, 940, 692, 194, 93, 397, 813, 577, 352, 439, 589, 967, 771, 938, 65, 973, 109, 913, 271, 514]. III [552, 440]. images [716]. imbedding [52, 474]. immersion [47]. Immersions [432, 892, 423]. imply [715]. incidence [12, 338, 438, 439, 440, 666]. inclined [165]. inclusion [762]. incomplete [16, 217, 321, 761, 348, 253, 1012]. increasing [79]. indecomposable [642]. indefinite [755]. indépendants [871]. independent [871, 372, 503, 702]. indeterminates [362]. index [410, 305, 219]. indice [410]. Indices [980]. induce [766]. Induced [319, 111, 139]. Inductive [810, 735, 690]. Inequalities [496, 897, 398, 865, 176, 629, 92]. inequality [971, 545, 64, 817]. Inertial [614]. Infinite [521, 337, 948, 285, 942, 403, 237, 507]. infinitely [65]. infinitesimal [358]. infinity [928, 946]. Ingham [450]. injective [776, 54]. injectivity [528, 480]. inner [858]. inscribed [613, 334]. inseparable [1010]. integer [438, 439, 440, 209]. integers [502, 845]. Integral [229, 12, 249, 258, 340, 607, 876, 377, 692, 498, 777, 108, 152, 720, 181, 318, 785, 351, 700, 472, 772, 201, 250, 802, 5, 67, 752, 887, 973, 529]. intégrale [108]. integrals [583, 574, 66, 539, 231]. Integration [87, 181, 172, 445]. interchanges [418]. interpolating [903]. interpolation [980, 570]. interpolatory [555]. Intersection [53]. intersections [268, 451]. Interval [257, 361, 314, 783]. intervals [617]. Intrinsic [233, 620, 855]. invariance [588, 763, 753]. Invariant [400, 464, 499, 479, 621, 643, 972, 919, 32, 710, 296]. Invariants [56, 817, 788, 449, 872]. inverse [335, 710, 694]. Inversion [487, 112, 35]. inversive [672]. involution [651, 733]. involutions [422, 878, 842]. involutorischer [651]. involving [398, 865, 947, 142]. irrationalities [236]. Irreducibility [170]. irreducible [362, 53, 836, 949]. isols [569]. Isometric [892, 419]. isometries [327, 799, 529]. isomorphic [985, 154]. isomorphism [568, 742, 887, 973]. Isomorphisms [800, 234, 419, 614, 808, 551]. isomorphs [667]. isomorphy [494]. isotopes [453]. Isotopy [459, 316, 77, 494]. isotopy-isomorphy [494]. Isotropic [423]. IV [108].

J. [785, 849]. Jacobi [426]. Jordan [44, 467].

Kähler [423]. Kakeya [427]. kernel [359, 505]. kernels [117, 484]. kind [595]. Kolmogorov [616]. konstanter [671]. Körpererweiterungen [965]. Kranzprodukten [754]. Kronecker [906]. Krull [812]. Krümmung [671]. Kulikov [975]. Kummer [655, 258]. Kurosh [364].

labelled [48, 82, 36, 162]. Lambek [191]. Lambert [2]. Laplace [475, 567, 487]. large [104]. latin [95, 17, 18, 660]. Lattice [381, 177, 26, 299, 538, 643, 954, 447, 86, 85, 163, 78, 929, 158, 769]. Lattice-ordered [381, 954]. lattices [425, 851, 698, 939, 940, 1011, 404, 826, 482, 272, 212, 720, 52, 597, 556, 166, 161, 38, 765, 798, 949, 341, 379, 872]. law [365]. Lebesgue [700, 88, 185]. left [343, 551]. Legendre [561, 580, 644]. lemma [919, 664, 315]. lemmas [903]. length [190, 344]. Levy [182]. Lie [1014, 169, 835, 721, 29, 757, 62, 143, 1001, 145, 144, 45, 81]. Lifting [826, 690]. like [132, 245, 584, 591, 322, 819]. limit [951]. Limits [38, 149, 235, 735, 690]. Lindelöf [649]. line [260, 406, 787, 781]. linéaires [108]. Linear [588, 763, 290, 313, 888, 548, 72, 844, 643, 190, 399, 498, 108, 476, 286, 212, 836, 34, 639, 114, 87, 854, 866, 916, 996, 563, 209, 150, 799, 909, 146, 665, 726, 55, 957, 958, 108]. Linearization [829]. Linearly [1000]. lines [587, 507, 517]. links [711]. Liouville [302, 67]. Lipschitz [61, 799]. Livingston [374]. Local [488, 58, 854, 802, 827, 529]. localement [796]. locally [753, 1011, 631, 953, 397, 522]. location [405]. loci [452]. logarithmic [822]. Longest [79]. look [655]. loop [453, 843, 896]. loop-isotopes [453]. loop-spaces [843, 896]. Loops [13, 266, 393, 494]. Lorentz [646]. Lototsky [303, 461, 156, 448, 292]. Lototsky-type [461]. low [225, 759]. low-order [225]. Lower [893, 629, 894, 734, 378, 933, 839, 920]. Lyapunov [416].

M. [792]. machine [477]. Mackey [1009]. Magic [541]. Majorants [445]. Malcev [191]. manifold [595, 488, 604]. manifolds [859, 860, 615, 892, 358, 591, 47, 918, 46, 208]. manuscript [735]. many [65, 925]. map [766]. mapping [547]. mappings [245, 477, 622]. Maps [766, 918, 738, 241, 317, 355, 544, 631, 630, 375, 443, 578, 159, 743, 214]. Markov [24]. martingale [482]. matrices [434, 588, 763, 496, 454, 706, 190, 693, 233, 368, 84, 164, 227, 659, 12, 207, 719, 438, 439, 440, 823, 307, 267, 240, 137, 223, 593, 69, 788, 41, 329, 470, 185, 39, 211, 709, 270, 418, 221]. Matrix [338, 761, 711, 94, 987, 588, 763, 890, 328, 1, 153, 579, 983, 275, 290, 23, 897, 70, 951, 520, 389, 300, 513, 123]. matroids [566, 988]. Maxima [390]. Maximal [716, 190, 18, 532, 944, 516, 305, 915, 255, 25, 746]. maximum [328, 869, 1013]. Mean [951, 950, 549]. means [101, 762, 60, 710]. measurability [247]. measurable [58, 884, 636]. measure [615, 810, 718, 830, 146]. Measures [514, 386, 777, 637, 782, 32]. meets [925]. meromorphic [634, 635, 235, 376, 808, 873]. metacyclic [978]. métaprimitifs [356]. Metaprimitive [356]. method [416, 822, 450, 388]. methods [926, 1, 153, 993, 88, 780, 956, 70]. metric [432, 533, 863, 755]. metrical [57]. metrics [572]. metrizability [768, 681, 301]. metrizable [828]. metrization [332, 512, 15, 188, 557]. Minimal [952, 546, 207, 418, 812, 164]. minimaux [871]. Minimum [147, 871, 187]. Minkowski [101]. minors [496]. Mixed [683, 550]. mod [921]. models [680]. Modifications [46]. Modular [492, 278, 425, 851, 1009, 933, 967, 872]. module [532]. modules [848, 111, 847, 436, 519, 576, 577, 530, 201, 776, 857]. modulo [789]. Moebius [373, 736]. moment [783, 102]. Monotone [212, 712, 900, 897, 622]. Montgomery [1006, 859, 860]. Moore [332, 501, 301]. Morita [990]. mosaics [431]. Motions [275, 37]. Moulton [99, 264, 323]. mountain [524]. mountain-climbing [524]. moyenne [950]. multiple [353, 752]. Multiplication [798, 363, 640, 633, 304]. Multiplicative [511]. Multiplicities [164, 555]. multiplier [800, 335, 298]. Multipliers [199, 391]. multivariate [723]. mutually [17].

Napierian [671]. Natural [745]. nature [300]. near [687, 994, 946]. near-rings [994]. Nearly [612]. negative [496, 706, 368, 711, 905, 69]. Neperschen [671]. net [504]. nets [504]. networks [410]. Neumann [735, 561]. Newman [792]. Next [555]. Next-to-interpolatory [555]. Nil [333, 469, 941]. nilpotency [896]. nilpotent [8, 29, 941, 40, 109]. nilpotents [29]. no [948, 435]. Nodal [44]. nodes [36, 162]. Noetherian [354, 110]. Non [985, 804, 253, 1012, 827, 257, 241, 317, 626, 496, 706, 20, 194, 368, 53, 705, 584, 820, 836, 289, 873, 59, 44, 863, 465, 201, 676, 711, 345, 767, 69, 726, 856, 819, 179]. non-associative [345, 767]. non-commutative [53, 44]. non-compact [873]. non-coprime [820]. non-decreasing [257]. non-degenerate [863, 819]. non-differentiable [705]. non-distributive [856]. non-Euclidean [676]. Non-existence [253, 1012, 626]. non-hyperfinite [985]. non-integral-domains [201]. Non-isomorphic [985]. non-linear [836, 726]. Non-local [827]. non-negative [496, 706, 368, 711, 69]. non-orientable [179]. non-planar [584]. non-residues [289]. non-self-adjoint [59]. non-separable [241, 317]. Non-stationary [804]. non-tangential [465]. nonabelian [20, 194]. nonlinear [523]. Normal [862, 106, 465, 434, 435, 404, 640, 12, 878, 811, 624, 211, 123]. normalizers [811]. normed [120, 121, 639]. norms [120]. Note [225, 531, 501, 500, 111, 605, 953, 266, 725, 963, 772, 586, 461, 842, 415, 27, 69, 909, 856, 211, 600, 943, 950]. Notes [1011, 585, 604]. null [24]. null-recurrent [24]. number [72, 463, 647, 220, 217, 18, 722, 518, 534, 328, 261, 256, 601, 824, 326, 82, 842, 660, 443, 732, 36, 383, 793, 197]. numbers [236, 877, 322, 744, 433]. numerators [537]. numerical [817].

observables [815]. octaèdre [410].

octahedra [26]. octahedron [410]. odd [96, 50]. odd-vertex-pairings [50]. Ohm [785]. One [294, 719, 110, 711, 325, 918, 925]. One-dimensional [294, 325]. one-parameter [925]. ones [328, 41]. only [19]. onto [159]. open [812, 76, 596]. operations [108, 255, 108]. operator [80, 972, 625, 302, 838, 820, 830, 437, 349]. Operators [825, 646, 618, 407, 511, 28, 151, 634, 635, 435, 364, 1007, 286, 756, 394, 59, 91, 106, 893, 916, 755, 203, 638, 730, 999, 909, 901]. orbits [9]. order [96, 115, 987, 28, 18, 248, 749, 225, 658, 893, 291, 866, 609, 885, 170, 905, 150, 389, 663, 751]. ordered [71, 381, 571, 954, 272, 490, 228, 694, 857]. ordering [625]. orders [745, 11, 949]. ordinal [877]. ordinary [618, 151, 687, 760, 104, 189]. ordinatizing [342]. ordonnés [571]. orientable [179]. orientations [50]. origin [946]. Orlicz [135]. Orr [4]. Orthogonal [659, 723, 17, 18, 816, 269, 95, 526, 802, 329]. orthogonality [666]. orthomodular [342]. Orthomorphisms [95]. orthonormal [90]. Oscillation [579, 983]. Osculating [180, 663]. other [268]. outwardly [260]. ovals [1003].

P. [849]. packed [287]. packing [636]. packings [784, 324, 585, 521]. pair [899]. paired [83]. pairings [50]. Pairs [289, 205, 491, 628, 1009]. pairwise [18]. paper [153]. Parabolic [242, 291, 176, 996, 649, 629, 92, 523]. paracompact [627, 885, 743, 301]. paracompactness [332]. Parallel [517, 146]. parameter [925]. part [498, 971]. Partial [1009, 865, 399, 380, 745, 405, 756, 563, 774]. partially [71, 625, 272, 348, 253, 1012]. Partition [96, 877, 695]. partitions [599, 757, 1014]. Pascal [1003]. pathology [379]. Paths [382, 237, 383]. patterns [463]. perfect [890, 225, 874, 456]. periodic [426, 330]. periodicity [509, 813]. Peripheral [715]. Permanent [137, 136, 513]. permanental [496, 545]. permis [410]. permitted [410]. Permutability [597]. permutation [384, 960]. permutations [328, 336]. Perron [426, 340, 607, 933]. person [505]. perspectivities [653]. Perturbation [28, 151]. perturbations [437, 523]. phenomenon [60, 185]. Phragmèn [649]. Pick [919]. Piecewise [550]. planar [241, 317, 584, 125, 214, 902]. plane [19, 132, 982, 658, 989, 406, 676, 598]. planes [831, 851, 284, 306, 879, 995, 369, 837, 265, 352, 963, 424, 99, 264, 323, 430, 1003, 297]. Poincaré [592, 312]. Point [245, 71, 48, 592, 900, 818, 961, 565]. Point-like [245]. points [281, 103, 147, 687, 590, 779, 820, 586, 842, 209, 596, 799, 665, 622]. pointwise [332, 301]. Polar [101]. polarities [369, 662, 665]. Pólya [641]. polygons [613, 75, 664, 155]. polyhedra [674, 256, 458, 402, 729, 771, 481]. polyhedral [510, 320, 819]. polynomial [120, 733, 747, 1005]. Polynomials [870, 880, 362, 816, 269, 490, 561, 448, 755, 170, 686, 570, 628, 224]. polytope [261, 326, 565, 346]. Polytopes [832, 677, 1000, 678]. Positive [803, 971, 823, 470, 986, 221, 300]. pour [410]. Power [3, 96, 42, 554, 282, 493, 821, 700, 289, 205, 15]. power-sets [15]. powers [948, 368, 302, 322, 349]. Prange [501]. preassigned [204]. prescribed [281, 377, 692, 376, 711, 870, 709]. presentation [834, 875]. presentations [661, 740]. presented [174]. preservers [137]. preserving [272]. Primary [554, 540, 890, 531]. Prime [436, 166, 72, 96, 42, 554, 498, 722, 519, 493, 838, 821, 305, 97, 957]. prime-power [42, 493, 821]. primes [619, 812]. primitive [960, 935, 750]. primitives [68, 750]. Principal [453, 218, 53, 526]. principle [974, 547]. probability [93, 503]. probalistic [713]. problem [548, 613, 360, 103, 244, 1009, 34, 82, 311, 387, 711, 849, 124, 636, 920, 667, 668, 524]. Problems [718, 115, 550, 756, 89, 31, 648, 783, 292, 370, 102, 168, 752, 726, 523, 827]. processes [804, 723, 680]. product [118, 454, 777, 429, 739, 249, 915, 637, 841, 858]. Products [719, 538, 149, 508, 527, 801, 639, 903, 543, 910, 729, 771, 803, 40, 109, 446]. Programmes [83]. projection [526]. Projections [154, 989]. Projective [847, 786, 325, 959, 553, 848, 284, 530, 369, 837, 850, 51, 265, 352, 963, 787, 598, 323, 133, 348, 430, 1003, 663, 690, 219]. projectives [283]. Proof [592, 616, 390, 678, 188]. proofs [512]. Properties [90, 203, 92, 697, 952, 444, 882, 605, 353, 316, 715, 77, 161, 823, 110, 488, 414]. property [734, 544, 58, 344, 494, 334]. Prüfer [540]. pseudo [582, 809, 463, 679]. pseudo-Euclidean [679]. pseudo-subordinate [582, 809]. pseudo-symmetric [463]. pseudometrics [536]. Pseudovaluations [686]. Ptolemaic [371]. pure [815, 43]. purely [1010].

quadrangles [265, 352]. quadrangular [367]. quadratic [229, 977, 529, 219]. quadrature [401]. quadrics [133]. quadruple [14]. quasi [548, 796, 466, 998, 776, 681, 913]. quasi-compacts [796]. quasi-fields [998]. quasi-framal [466]. quasi-Frobenius [913]. quasi-injective [776]. quasi-linear [548]. quasi-metrizability [681]. quasigroups [453]. quaternions [693, 233]. quelques [2]. question [981]. questions [812]. quota [359]. quotients [381, 937, 97, 226, 255, 573].

R. [792]. Radial [235]. radical [734, 701, 605, 116, 10]. radicals [396, 881, 116, 364, 894]. Ramanujan [699, 868, 898, 142]. random [616, 503]. Randomly [287]. range [616, 542, 933]. range-correspondences [542]. rank [454, 207, 408, 869, 1013]. rate [651]. rates [503]. Rational [654, 209, 882, 532, 338, 3, 433]. Räumen [671]. Real [118, 233, 89, 971, 203, 7]. reality [121]. realizable [425]. rearrangement [621]. rearrangement-invariant [621]. rearrangements [673]. Reciprocal [66]. reciprocals [236]. Reconstruction [992, 516]. rectangles [660]. Rectangular [439, 440]. Recurrent [907, 24]. recurring [190]. Redfield [215, 542]. reduced [660]. reducibility [285, 167]. reduction [215, 711]. Refinement [361]. Refinement-unbounded [361]. refinements [416]. reflection [974, 56]. reflections [98]. reflexive [911]. Regel [671]. regional [518]. regions [78]. regressive [589]. Regular [674, 408, 738, 546, 626, 75, 1, 153, 735, 612, 458, 691, 915, 88, 366, 240, 185, 471, 795, 925, 54, 480, 943]. regularity [474, 853]. Reidemeister [837]. related [236, 82]. relating [946]. relation [832]. Relations [190, 322, 290, 228, 877, 596, 875]. relationship [980, 470]. relationships [399]. relative [190]. relatively [838, 836, 32, 765]. remark [530]. remarkable [222]. Remarks [447, 303, 818, 993, 661, 740]. Remarque [950]. Remetrization [927]. remplissage [891]. Rényi [616]. Replaceable [504]. Representation [640, 209, 777, 108, 451, 502, 392, 765, 108]. Representations [144, 738, 721, 29, 139, 11, 288, 249, 145, 319, 846, 278, 457, 803, 29]. represented [713]. representing [3]. réseaux [410]. residue [217]. residues [289, 611, 205, 628]. respect [583, 760]. restricted [462, 522]. Restrictive [807]. result [173, 792]. results [17, 876, 156, 303, 616, 633, 503, 206, 946, 133]. revolution [657]. Ricci [315]. Riemann [700, 434, 464, 873, 67, 479]. Riemann-type [700]. Riemannian [671, 196, 604]. Riemannschen [671]. Riesz [609, 762]. Right [998, 218]. rigid [558, 934]. ring [906, 852, 252, 708, 365, 226, 304, 255, 223, 10, 480, 551, 708]. ringed [984]. Rings [733, 937, 873, 773, 778, 396, 381, 96, 246, 354, 997, 363, 701, 890, 568, 742, 605, 116, 364, 633, 126, 408, 735, 614, 193, 333, 469, 218, 275, 854, 97, 941, 994, 685, 22, 874, 776, 350, 887, 973, 471, 1002, 313, 356, 54, 894, 234, 263]. robin [428]. root [746]. rooted [317, 375, 578]. roots [935, 570]. rotation [759, 1004]. rotations [219]. round [428]. row [711]. rule [671].

Samelson [1006, 859, 860]. satisfy [120, 121]. satisfying [12, 333, 469, 338, 778, 315]. saturated [409]. Schauder [712, 357, 564]. Schlicht [370]. Schmidt [544]. Scholz [920]. Schwarz [919]. Schwarzian [866]. Screenability [332]. second [416, 150, 549]. second-order [150]. sections [989]. Seidel [981]. selection [884]. Self [535, 59, 755, 999, 54, 480]. self-adjoint [755, 999]. Self-centered [535]. self-injectivity [480]. Semi [631, 519, 724, 908, 582, 809, 483, 620, 995, 97, 62, 750, 310, 819, 54]. semi-algebras [582, 809]. semi-continuous [819]. Semi-groups [631, 62]. Semi-homomorphisms [724]. Semi-prime [519, 97]. semi-simple [483, 620, 310, 54]. semi-simples [750]. semi-translation [995]. Semi-valuations [908]. semidefinite [300]. semigroup [556, 661, 740]. Semigroups [342, 343, 485, 807, 805, 638, 694, 602, 741, 913]. semilattice [597]. semimodular [52]. Semisimple [750]. semisimplicity [967]. separable [241, 317, 848, 847]. separate [118]. separation [279, 853]. sequence [923, 502, 714, 473, 564]. sequences [90, 617, 190, 186, 227, 330, 415, 468, 632, 806]. Series [515, 569, 68, 882, 861, 282, 63, 442, 378, 66, 561, 3, 905, 904, 478, 644, 2, 2]. set [339, 268, 103, 962, 451, 760, 863, 839]. set-functions [339]. Sets [816, 165, 587, 71, 135, 1006, 280, 871, 706, 190, 268, 198, 230, 826, 272, 58, 476, 749, 845, 175, 989, 815, 758, 335, 814, 535, 581, 391, 525, 589, 555, 199, 788, 517, 636, 146, 961, 15, 202, 918, 298]. seven [958]. several [362]. sextics [982]. Sheets [7]. Sieve [617, 468]. Sieve-generated [617, 468]. Signal [572]. Silov [563]. similitudes [127, 128]. Simple [74, 467, 19, 721, 260, 483, 620, 650, 310, 141, 54, 895, 234]. simples [750]. Simplicial [481, 245]. simplicity [256]. simply [960]. Simultaneous [239, 788, 772]. simultaneously [816]. Singular [663, 523, 1006, 339, 618, 687, 91, 210, 833]. singularities [405]. size [731]. skew [987, 674, 307]. skew-symmetric [307]. Slepian [501]. slicings [184]. small [503, 862]. small-deviation [503]. Smirnov [616]. smoothness [484]. Sobolev [544, 474]. solidly [287]. Soluble [652, 969]. solution [318, 1009, 311, 387, 331, 171]. solutions [129, 217, 550, 760, 176, 563, 629, 92, 105, 414]. Solvability [687, 293]. solvable [422]. Some [416, 399, 156, 303, 98, 812, 353, 175, 633, 818, 503, 884, 840, 89, 903, 993, 822, 143, 916, 324, 783, 441, 110, 623, 240, 946, 49, 661, 740, 142, 133, 348, 379, 913, 243, 398, 616, 419, 656, 225, 254, 780, 956, 897, 253, 1012, 726, 2, 414]. sommets [871]. Sonnenschein [185]. Space [208, 177, 553, 646, 407, 790, 108, 631, 152, 974, 674, 756, 106, 6, 324, 609, 761, 137, 223, 506, 799, 909, 146, 663, 901, 105, 499]. space-time [506]. space-valued [974]. spaces [432, 239, 135, 119, 21, 671, 768, 621, 980, 118, 544, 283, 147, 154, 476, 533, 562, 779, 911, 316, 332, 758, 843, 896, 77, 828, 639, 850, 914, 83, 228, 863, 916, 755, 934, 627, 624, 976, 885, 623, 70, 858, 348, 295, 564, 645, 180, 663, 743, 984, 795, 301, 927, 888, 417, 179, 414, 557, 243]. Special [814, 227, 22]. spectral [923, 777, 714, 823]. spectrum [753, 28, 151, 893, 804]. sphere [245, 601, 324, 585, 731]. spheres [177, 953, 287, 601, 784, 165, 478, 668]. spherical [953, 704]. spherically [171]. spin [392]. spiral [981, 606]. Splitting [589]. Square [438, 305]. squares [17, 18, 225, 95, 192, 209]. Stability [760, 104]. stable [797, 1000]. Standard [721]. star [819]. star-like [819]. starlike [168]. Stasheff [867]. states [815]. static [171]. stationary [804, 723]. Steiner [565]. Stieltjes [487, 472]. stochastic [137, 723, 470, 513]. Stone [939, 940]. Stonian [223]. strength [329]. strong [182]. strongly [931, 764, 927, 943]. Structure [859, 860, 940, 485, 178, 395, 368, 840, 97]. structures [774, 666]. Study [128, 790]. Sturm [302]. sub [499]. sub-bundles [499]. subalgebras [725, 158]. subdivision [449]. subduced [144]. subgraph [1008]. subgraphs [952, 702]. subgroup [42, 797]. subgroups [889, 978, 422, 193, 305, 446]. Sublattices [85, 86, 163, 856]. Submethods [1, 153]. submodules [43]. Subnormal [193]. Subordinate [582, 809]. subplanes [430, 297]. Subrings [255, 773, 941, 25]. subsemigroup [925]. subsemigroups [343, 787]. subsequences [79]. subset [444]. subsets [1005]. subspace [643, 464, 196]. subspaces [972, 154, 400, 87, 357, 209, 712, 146, 479]. subsystems [746]. subtrees [516]. successive [149]. sum [236, 209, 505]. Summability [330, 561, 70, 926, 186, 1, 153, 728, 993, 822, 905, 560, 388]. Summability- [330]. summable [442]. summation [947, 450]. Sums [274, 33, 537, 433, 865, 890, 749, 502, 801, 503, 711, 142, 322, 774]. Supermagic [603]. surface [655]. surfaces [654, 464, 148, 420, 873, 479, 179]. suspension [921]. suspensions [843]. Sylow [80, 157, 978, 422]. Symmetric [645, 130, 463, 147, 911, 307, 392, 846, 677, 803, 171]. symmetries [527, 216]. symmetry [386]. system [563]. Systems [932, 151, 399, 217, 14, 104, 914, 786, 262, 976, 665, 65, 746]. Szego [880, 490].

tables [640]. Tabling [703]. tactical [254]. tangent [499]. tangential [149, 465]. Tauberian [926, 722, 231, 822, 366]. Taylor [882, 60]. Tchebycheff [401]. temperature [515]. tensor [739, 915, 803]. tensors [790, 514]. Term [454, 684, 207, 869, 1013]. terms [502]. tetrahedra [299]. Tetraspheres [100]. th [866]. their [980, 149, 982, 97]. theorem [71, 571, 486, 926, 880, 200, 374, 921, 355, 991, 722, 482, 101, 900, 409, 351, 757, 1014, 312, 43, 762, 390, 765, 593, 675, 182, 57, 501, 641, 678, 446, 298, 857, 895]. théorème [571]. theorems [71, 72, 592, 115, 191, 259, 939, 940, 487, 431, 464, 231, 884, 474, 822, 49, 951, 649, 188]. theory [859, 860, 906, 80, 738, 280, 93, 656, 172, 472, 757, 1014, 134, 786, 311, 387, 752, 65, 912, 457, 178, 988]. thickness [421]. thirteen [668]. Three [982, 98, 674, 488, 203, 751, 875]. three-manifold [488]. three-space [674]. Time [703, 506]. Time-Tabling [703]. Topological [420, 828, 768, 1011, 508, 316, 77, 417, 263]. Topologies [538, 858, 608, 447, 769]. topology [1008]. Torsion [201]. Torsion-free [201]. total [346]. Totally [758]. Tournaments [309, 428]. Trace [169, 302]. Traces [267, 41]. tracts [543]. Transcendental [126]. transform [475, 112, 621, 156, 303, 487, 448, 560]. Transformation [113, 763, 643, 907, 945, 87, 966]. Transformations [962, 72, 844, 588, 567, 244, 269, 114, 358, 373, 736, 996, 761, 366, 461, 888]. transforms [292, 946, 580]. transitifs [336]. transitive [960, 306, 559]. transitivity [384]. translated [617]. translation [995, 61, 341]. Transversal [988]. transverse [968]. tree [516, 578, 411]. tree-equivalent [411]. tree-rooted [578]. Trees [510, 382, 411, 48, 244, 794, 786, 443]. triangle [599, 334]. Triangles [189, 222]. triangular [375]. Triangulation [610]. triangulations [125]. trigonometric [68, 102]. trigonometrical [63]. triples [647, 694]. trivalent [402]. Turán [390, 849]. twelve [546]. Two [876, 684, 431, 854, 667, 697, 672, 98, 344, 31, 996, 525, 292, 237, 517, 329, 202, 551]. Two-dimensional [854, 98]. two-way [237]. type [4, 926, 42, 552, 626, 616, 720, 351, 539, 700, 330, 74, 461, 25, 959, 716]. types [713]. typically [89].

Ulam [486]. ultra [704]. ultra-spherical [704]. unary [970]. Unbounded [986, 361, 1010]. Undirected [425, 558]. unequal [784]. unified [171]. Uniform [119, 284, 140, 357, 557]. uniformities [917]. Union [196, 751]. Unions [202, 525]. Unique [637, 1002]. Uniqueness [495]. unit [370, 168, 731]. unitaires [29]. Unitary [846, 98, 29, 223, 788, 457, 39, 123]. unity [570]. Univalent [129, 606, 950]. univalentes [950]. universal [638]. unodes [654]. unsolvable [619]. unsymmetrical [253, 1012]. upper [819]. Upsets [428]. use [770]. Utumi [226].

valence [801]. valuation [997, 786]. valuations [908, 832]. value [548, 115, 550, 951, 523, 827]. valued [974, 223, 247]. values [204, 549, 1005]. vanishing [928, 886]. variables [697, 616, 498, 405, 503, 203]. variant [758, 668]. Variation [269, 230, 583, 308]. Variation-diminishing [269]. variational [445, 523]. variations [752]. varieties [553, 7]. Variety [872]. Vector [720, 358, 698, 826, 482, 810, 758, 166, 798, 247, 379]. vector-valued [247]. vectors [823]. Verallgemeinerung [671]. vertex [50]. vertices [871, 326]. very [660]. VI [29, 663]. Volterra [472]. volume [919].

way [237]. Weak [139, 279, 924, 171, 414]. Weakly [135, 483, 723, 221]. wedge [200]. Weierstrass [112, 586]. weight [457]. well [415, 806]. well-distributed [415, 806]. which [121, 790, 107, 344, 9, 422, 601, 766, 343, 773, 710, 925]. Whittaker [203]. whose [120, 948, 878, 819]. Width [227]. Widths [84, 164]. Wiener [574, 182]. word [311, 387].

Zero [543, 195, 659, 766, 719, 145, 294]. zeroone [719]. zeros [876, 328, 5, 41]. Zur [754, 965, 671].

References

Goffman:1956:SRM

[1] Casper Goffman and G. M. Petersen. Submethods of regular matrix summability methods. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):40–46, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [153].

Touchard:1960:QSL

[2] Jacques Touchard. Sur quelques séries de Lambert et de Dirichlet. (French) [On some series of Lambert and Dirichlet]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):1–19, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Melzak:1960:PSR

[3] Z. A. Melzak. Power series representing certain rational functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 20–26, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Agarwal:1960:CBH

[4] Nirmala Agarwal. Certain bilateral hypergeometric identities of Cayley and Orr type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):27–39, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:1960:CBG

[5] Q. I. Rahman. On connections between growth and distribution of zeros of integral functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):40–44, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kurepa:1960:CFE

[6] Svetozar Kurepa. A cosine functional equation in Hilbert space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 45–50, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wallace:1960:SRA

[7] Andrew H. Wallace. Sheets of real analytic varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):51–67, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bachman:1960:FNG

[8] G. Bachman. On finite nilpotent groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):68–72, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gorenstein:1960:FGW

[9] Daniel Gorenstein. Finite groups which admit an automorphism with few orbits. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):73–100, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thierrin:1960:RCA

[10] G. Thierrin. Sur le radical corpoïdal d’un anneau. (French) [On the corpoidal radical of a ring]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 101–106, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Higman:1960:ROD

[11] D. G. Higman. On representations of orders over Dedekind domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 107–125, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldhaber:1960:IAN

[12] J. K. Goldhaber. Integral p-adic normal matrices satisfying the incidence equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):126–133, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cowell:1960:LA

[13] W. R. Cowell. Loops with adjoints. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12 (??):134–144, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hanani:1960:QS

[14] Haim Hanani. On quadruple systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12 (??):145–157, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Silverman:1960:MPS

[15] Robert Silverman. A metrization for power-sets with applications to combinatorial analysis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):158–176, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bose:1960:CBI

[16] R. C. Bose and S. S. Shrikhande. On the composition of balanced incomplete block designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):177–188, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bose:1960:FRC

[17] R. C. Bose, S. S. Shrikhande, and E. T. Parker. Further results on the construction of mutually orthogonal Latin squares and the falsity of Euler’s conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):189–203, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chowla:1960:MNP

[18] S. Chowla, P. Erdős, and E. G. Straus. On the maximal number of pairwise orthogonal Latin squares of a given order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12 (??):204–208, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bing:1960:SCC

[19] R. H. Bing. A simple closed curve is the only homogeneous bounded plane continuum that contains an arc. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 209–230, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dedecker:1960:CNA

[20] Paul Dedecker. Sur la cohomologie non abélienne. I. (French) [On nonabelian cohomology. I]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):231–251, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banaschewski:1960:HBE

[21] Bernhard Banaschewski. On homeomorphisms between extension spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12 (??):252–262, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martindale:1960:CSC

[22] Wallace S. Martindale III. The commutativity of a special class of rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12 (??):263–268, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1960:MC

[23] Marvin Marcus and Nisar A. Khan. On matrix commutators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 269–277, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lamperti:1960:NRM

[24] John Lamperti. On null-recurrent Markov chains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):278–288, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pukanszky:1960:MAS

[25] Lajos Pukánszky. On maximal albelian subrings of factors of type II_{1}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 289–296, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mordell:1960:LO

[26] L. J. Mordell. Lattice octahedra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 297–302, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Oppenheim:1960:NCF

[27] A. Oppenheim. A note on continued fractions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):303–308, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butler:1960:PCS

[28] John B. Butler, Jr. Perturbation of the continuous spectrum of even order differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):309–323, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dixmier:1960:RUG

[29] Jacques Dixmier. Sur les représentations unitaires des groupes de Lie nilpotents. VI. (French) [On the unitary representations of nilpotent Lie groups. VI]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12 (??):324–352, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1960:CFA

[30] F. W. Anderson. A class of function algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):353–362, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kiefer:1960:ETE

[31] J. Kiefer and J. Wolfowitz. The equivalence of two extremum problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12 (??):363–366, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mahowald:1960:RIM

[32] Mark Mahowald. On relatively invariant measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):367–373, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stewart:1960:SFD

[33] B. M. Stewart. Sums of functions of digits. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):374–389, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1960:LDP

[34] S. M. Johnson. A linear diophantine problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):390–398, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:1960:AIF

[35] Eckford Cohen. Arithmetical inversion formulas. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):399–409, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Read:1960:NCG

[36] R. C. Read. The number of k-colored graphs on labelled nodes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 410–414, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Greenberg:1960:DGM

[37] Leon Greenberg. Discrete groups of motions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):415–426, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Macbeath:1960:LLC

[38] A. M. Macbeath and S. Świerczkowski. Limits of lattices in a compactly generated group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):427–437, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stander:1960:CFC

[39] J. W. Stander and N. A. Wiegmann. Canonical forms for certain matrices under unitary congruence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 438–446, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Struik:1960:NPC

[40] Ruth Rebekka Struik. On nilpotent products of cyclic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 447–462, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ryser:1960:TMZ

[41] H. J. Ryser. Traces of matrices of zeros and ones. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):463–476, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brahana:1960:NTC

[42] H. R. Brahana. A new type of characteristic subgroup of primepower groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):477–482, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kolettis:1960:TPS

[43] George Kolettis, Jr. A theorem on pure submodules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):483–487, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kokoris:1960:NNC

[44] Louis A. Kokoris. Nodal noncommutative Jordan algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 488–492, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ree:1960:GLE

[45] Rimhak Ree. Generalized Lie elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12 (??):493–502, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wallace:1960:MCM

[46] Andrew H. Wallace. Modifications and cobounding manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):

503–528, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Samelson:1960:IM

[47] Hans Samelson. On immersion of manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):529–534, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Austin:1960:EPL

[48] T. L. Austin. The enumeration of point labelled chromatic graphs and trees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12 (??):535–545, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nettleton:1960:SGT

[49] R. E. Nettleton. Some generalized theorems on connectivity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 546–554, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nash-Williams:1960:OCO

[50] C. St. J. A. Nash-Williams. On orientations, connectivity and odd-vertex-pairings in finite graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 555–567, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kelly:1960:PCC

[51] Paul Kelly and E. G. Straus. On the projective centres of convex curves.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 12 (??):568–581, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Finkbeiner:1960:SIL

[52] D. T. Finkbeiner. A semimodular imbedding of lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):582– 591, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feller:1960:III

[53] Edmund H. Feller. Intersection irreducible ideals of a non-commutative principal ideal domain. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):592– 596, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Utumi:1960:CRR

[54] Yuzo Utumi. On continuous regular rings and semi-simple self injective rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):597–605, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steinberg:1960:AFL

[55] Robert Steinberg. Automorphisms of finite linear groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):606–615, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steinberg:1960:IFR

[56] Robert Steinberg. Invariants of finite reflection groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):616–618, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schmidt:1960:MTD

[57] Wolfgang Schmidt. A metrical theorem in diophantine approximation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 619–631, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eames:1960:LPM

[58] W. Eames. A local property of measurable sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):632–640, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kemp:1960:CNS

[59] R. R. D. Kemp. On a class of non-self-adjoint differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 641–659, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miracle:1960:GPT

[60] Chester L. Miracle. The Gibbs phenomenon for Taylor means and for [F, d_{n}] means. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):660–673, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mirkil:1960:CTH

[61] H. Mirkil. Continuous translation of Hölder and Lipschitz functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??): 674–685, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Langlands:1960:LSG

[62] R. P. Langlands. On Lie semi-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12 (??):686–693, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cross:1960:ETS

[63] George Cross. The expression of trigonometrical series in Fourier form. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12 (??):694–698, ???? 1960. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mordell:1960:DI

[64] L. J. Mordell. On a discriminant inequality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 12(??):699–704, ???? 1960. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Segal:1961:FTD

[65] I. E. Segal. Foundations of the theory of dynamical systems of infinitely many degrees of freedom. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):1–18, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Guinand:1961:RCC

[66] A. P. Guinand. Reciprocal convergence classes for Fourier series and integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):19–36, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Riesz:1961:ACR

[67] Marcel Riesz. The analytic continuation of the Riemann–Liouville integral in the hyperbolic case. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):37–47, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bullen:1961:CPG

[68] P. S. Bullen. Construction of primitives of generalized derivatives with applications to trigonometric series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 48–58, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putnam:1961:NNN

[69] C. R. Putnam. A note on nonnegative matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):59–62, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mitchell:1961:SMM

[70] Josephine Mitchell. Summability methods on matrix spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):63–77, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Abian:1961:TPO

[71] Smbat Abian and Arthur B. Brown. A theorem on partially ordered sets, with applications to fixed point theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):78–82, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Amitsur:1961:ALT

[72] S. A. Amitsur. Arithmetic linear transformations and abstract prime number theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):83–109, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [844].

Hoffman:1961:BDG

[73] A. J. Hoffman and Moses Richardson. Block design games. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):110–128, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kleinfeld:1961:SAT

[74] Erwin Kleinfeld. Simple algebras of type (1, 1) are associative. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 129–148, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Crowe:1961:GRC

[75] D. W. Crowe. The groups of regular complex polygons. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):149–156, ????

1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nash-Williams:1961:DFG

[76] C. St. J. A. Nash-Williams. Decomposition of finite graphs into open chains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):157–166, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hu:1961:HIP

[77] Sze tsen Hu. Homotopy and isotopy properties of topological spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 167–176, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Macbeath:1961:CFR

[78] A. M. Macbeath. On convex fundamental regions for a lattice. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 177–178, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schensted:1961:LID

[79] C. Schensted. Longest increasing and decreasing subsequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 179–191, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ayoub:1961:STC

[80] Christine W. Ayoub. Sylow theory for a certain class of operator groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 192–200, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Togo:1961:DAL

[81] Shigeaki Tôgô. On the derivation algebras of Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):201–216, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:1961:EPR

[82] C. Y. Lee. An enumeration problem related to the number of labelled bicoloured graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):217–220, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kretschmer:1961:PPS

[83] K. S. Kretschmer. Programmes in paired spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):221–238, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fulkerson:1961:WHM

[84] D. R. Fulkerson and H. J. Ryser. Widths and heights of (0, 1)-matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):239–255, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jonsson:1961:SFL

[85] Bjarni Jónsson. Sublattices of a free lattice. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):256–264, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Galvin:1961:DSF

[86] Fred Galvin and Bjarni Jónsson. Distributive sublattices of a free lattice. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):265–272, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kobayashi:1961:ISD

[87] Edward T. Kobayashi. Integration of subspaces derived from a linear transformation field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):273–282, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorch:1961:LCR

[88] Lee Lorch and Donald J. Newman. The Lebesgue constants for regular Hausdorff methods. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):283–298, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jenkins:1961:SPT

[89] James A. Jenkins. Some problems for typically real functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 299–304, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bullen:1961:PCO

[90] P. S. Bullen. Properties of the coefficients of orthonormal sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 305–315, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kemp:1961:CSD

[91] R. R. D. Kemp. On a class of singular differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 316–330, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Protter:1961:PSP

[92] M. H. Protter. Properties of solutions of parabolic equations and inequalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):331–345, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdos:1961:GTP

[93] P. Erdős. Graph theory and probability. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):346–352, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smiley:1961:MC

[94] M. F. Smiley. Matrix commutators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):353–355, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1961:OGO

[95] Diane M. Johnson, A. L. Dulmage, and N. S. Mendelsohn. Orthomorphisms of groups and orthogonal latin squares. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):356–372, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Appel:1961:PRC

[96] K. I. Appel. Partition rings of cyclic groups of odd prime power order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):373–391, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lambek:1961:SSP

[97] Joachim Lambek. On the structure of semi-prime rings and their rings of quotients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):392–417, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Crowe:1961:STD

[98] D. W. Crowe. Some two-dimensional unitary groups generated by three reflections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):418–426, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pierce:1961:MP

[99] William A. Pierce. Moulton planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):427–436, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deMajo:1961:T

[100] A. de Majo. Tetraspheres. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 437–443, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Firey:1961:PMC

[101] William J. Firey. Polar means of convex bodies and a dual to the Brunn– Minkowski theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):444–453, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1961:HTM

[102] P. G. Rooney. On the Hausdorff and trigonometric moment problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):454–461, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cole:1961:EPB

[103] Randal H. Cole. The expansion problem with boundary conditions at a finite set of points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):462–479, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartman:1961:SLS

[104] Philip Hartman. On stability in the large for systems of ordinary differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):480–492, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Swanson:1961:ASE

[105] C. A. Swanson and M. Schulzer. Asymptotic solutions of equations in Banach space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):493–504, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Krabbe:1961:NOB

[106] Gregers L. Krabbe. Normal operators on the Banach space L^{p}(−∞, ∞). I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):505–518, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burgess:1961:HCW

[107] C. E. Burgess. Homogeneous continua which are almost chainable. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 519–528, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dinculeanu:1961:RIC

[108] N. Dinculeanu and C. Foias. Sur la représentation intégrale des certaines opérations linéaires. IV. opérations linéaires sur l’espace Ł_{a}^{p}. (French) [On the integral representation of certain linear operations. IV. Linear operations on the space Ł_{a}^{p}]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):529–556, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Struik:1961:NPC

[109] Ruth Rebekka Struik. On nilpotent products of cyclic groups. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 557–568, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Matlis:1961:SPN

[110] Eben Matlis. Some properties of Noetherian domains of dimension one. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13 (??):569–586, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Curtis:1961:NIM

[111] Charles W. Curtis. A note on induced modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):587–592, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bilodeau:1961:IFW

[112] G. G. Bilodeau. An inversion formula for the Weierstrass transform. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 593–601, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

LeBlanc:1961:TA

[113] Léon LeBlanc. Transformation algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):602–613, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kestelman:1961:ALT

[114] H. Kestelman. Anticommuting linear transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):614–624, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barrett:1961:FOB

[115] John H. Barrett. Fourth order boundary value problems and comparison theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):625–638, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Divinsky:1961:GRC

[116] N. Divinsky. General radicals that coincide with the classical radical on rings with D.C.C. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):639–644, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barros-Neto:1961:AK

[117] J. Barros-Neto and F. E. Browder. The analyticity of kernels. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 645–649, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Browder:1961:RAF

[118] Felix E. Browder. Real analytic functions on product spaces and separate analyticity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):650–656, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Atsuji:1961:UCC

[119] Masahiko Atsuji. Uniform continuity of continuous functions on uniform spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):657–663, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aurora:1961:NAWa

[120] Silvio Aurora. On normed algebras whose norms satisfy polynomial identities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):664–674, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aurora:1961:NAWb

[121] Silvio Aurora. On normed algebras which satisfy a reality condition. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??): 675–682, ???? 1961. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wright:1961:CCG

[122] E. M. Wright. Counting coloured graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):683–693, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Youla:1961:NFM

[123] D. C. Youla. A normal form for a matrix under the unitary congruence group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 13(??):694–704, ???? 1961. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Read:1962:CCG

[124] R. C. Read. Contributions to the cell growth problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):1–20, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1962:CPT

[125] W. T. Tutte. A census of planar triangulations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):21–38, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Halperin:1962:TEC

[126] Israel Halperin. Transcendental elements in continuous rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 39–44, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wonenburger:1962:CAG

[127] María J. Wonenburger. The Clifford algebra and the group of similitudes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14 (??):45–59, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wonenburger:1962:SCS

[128] María J. Wonenburger. Study of certain similitudes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):60–68, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1962:US

[129] R. K. Brown. Univalent solutions of W′′ + pW = 0. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):69–78, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shapiro:1962:SED

[130] Victor L. Shapiro. Symmetric exterior differentiation and flat forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 79–86, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vanstone:1962:GFG

[131] J. R. Vanstone. A generalization of Finsler geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):87–112, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bing:1962:ECL

[132] R. H. Bing. Embedding circle-like continua in the plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):113–128, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ray-Chaudhuri:1962:SRQ

[133] D. K. Ray-Chaudhuri. Some results on quadrics in finite projective geometry based on Galois fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 129–138, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kleisli:1962:HTA

[134] Heinrich Kleisli. Homotopy theory in abelian categories. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):139–169, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ando:1962:WCS

[135] Tsuyoshi Andô. Weakly compact sets in Orlicz spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):170–176, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1962:PF

[136] Marvin Marcus and F. C. May. The permanent function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):177–189, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moyls:1962:PPS

[137] B. N. Moyls, Marvin Marcus, and Henryk Minc. Permanent preservers on the space of doubly stochastic matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):190–194, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gorenstein:1962:FGF

[138] Daniel Gorenstein. On finite groups of the form ABA. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):195–236, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fell:1962:WCI

[139] J. M. G. Fell. Weak containment and induced representations of groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14 (??):237–268, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glicksberg:1962:UBG

[140] Irving Glicksberg. Uniform boundedness for groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):269–276, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steinberg:1962:GSG

[141] Robert Steinberg. Generators for simple groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):277–283, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nicol:1962:SFI

[142] C. A. Nicol. Some formulas involving Ramanujan sums. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):284–286, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Laufer:1962:SLA

[143] P. J. Laufer and M. L. Tomber. Some Lie admissible algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 287–292, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Noonan:1962:RSI

[144] B. Noonan. Representations subduced on an ideal of a Lie algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 293–303, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Noonan:1962:CRL

[145] B. Noonan. The construction of representations of Lie algebras of characteristic zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):304–312, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Santalo:1962:MSP

[146] L. A. Santaló. On the measure of sets of parallel linear subspaces in affine space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14 (??):313–319, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Crittenden:1962:MCP

[147] Richard J. Crittenden. Minimum and conjugate points in symmetric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14 (??):320–328, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Helfenstein:1962:AHS

[148] H. G. Helfenstein. Analytic Hopf surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):329–333, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cargo:1962:ATL

[149] G. T. Cargo. Angular and tangential limits of Blaschke products and their successive derivatives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 334–348, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Petty:1962:GAS

[150] C. M. Petty and J. E. Barry. A geometrical approach to the second-order linear differential equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 349–358, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butler:1962:PCS

[151] John B. Butler, Jr. Perturbation of the continuous spectrum of systems of ordinary differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 359–378, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eberlein:1962:IFS

[152] W. F. Eberlein. An integral over function space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):379–383, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goffman:1962:CPS

[153] Casper Goffman and G. M. Petersen. Correction to the paper: “Submethods of regular matrix summability methods”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):384, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [1].

Dean:1962:PCC

[154] David W. Dean. Projections in certain continuous function spaces C(H) and subspaces of C(H) isomorphic with C(H). Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):385–401, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1962:CHP

[155] W. T. Tutte. A census of Hamiltonian polygons. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):402–417, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cowling:1962:SRG

[156] V. F. Cowling and C. L. Miracle. Some results for the generalized Lototsky transform. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):418–435, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrections and remarks [303].

Brauer:1962:FGA

[157] Richard Brauer and Henry S. Leonard, Jr. On finite groups with an abelian Sylow group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):436–450, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sachs:1962:LSB

[158] David Sachs. The lattice of subalgebras of a Boolean algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 451–460, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Namioka:1962:CM

[159] Isaac Namioka. On certain onto maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14 (??):461–466, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sion:1962:ACD

[160] Maurice Sion. Approximate continuity and differentiation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):467–475, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jonsson:1962:APF

[161] Bjarni Jónsson. Arithmetic properties of freely α-generated lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 476–481, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Read:1962:EGL

[162] Ronald C. Read. Euler graphs on labelled nodes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):482–486, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jonsson:1962:FSF

[163] Bjarni Jónsson and James E. Kiefer. Finite sublattices of a free lattice. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14 (??):487–497, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fulkerson:1962:MMW

[164] D. R. Fulkerson and H. J. Ryser. Multiplicities and minimal widths for (0, 1)-matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):498–508, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mauldon:1962:SEI

[165] J. G. Mauldon. Sets of equally inclined spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):509–516, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1962:PIV

[166] D. G. Johnson and J. E. Kist. Prime ideals in vector lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 517–528, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1962:CRG

[167] Diane M. Johnson, A. L. Dulmage, and N. S. Mendelsohn. Connectivity and reducibility of graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):529–539, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Royster:1962:EPF

[168] W. C. Royster. Extremal problems for functions starlike in the exterior of the unit circle. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):540–552, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Block:1962:TFL

[169] Richard Block. Trace forms on Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):553–564, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCarthy:1962:IBP

[170] P. J. McCarthy. Irreducibility of Bernoulli polynomials of higher order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14 (??):565–567, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vanstone:1962:GSS

[171] J. R. Vanstone. The general static spherically symmetric solution of the “weak” unified field equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 568–576, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hayes:1962:TI

[172] Charles A. Hayes, Jr. A theory of integration. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):577–596, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kiefer:1962:ER

[173] J. Kiefer. An extremum result. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 597–601, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Macdonald:1962:CFP

[174] I. D. Macdonald. On a class of finitely presented groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):602–613, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gordon:1962:SND

[175] Basil Gordon, W. H. Mills, and L. R. Welch. Some new difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 614–625, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lees:1962:ABS

[176] Milton Lees. Asymptotic behaviour of solutions of parabolic differential inequalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):626–631, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bleicher:1962:LCS

[177] M. N. Bleicher. Lattice coverings of n-space by spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):632–650, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tomiuk:1962:STC

[178] Bohdan J. Tomiuk. Structure theory of complemented Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 651–659, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yang:1962:NOC

[179] C. T. Yang. On the non-orientable closed surfaces in Euclidean spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14 (??):660–668, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Scherk:1962:OS

[180] Peter Scherk. Osculating spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 669–684, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdelyi:1962:FID

[181] A. Erdélyi and I. N. Sneddon. Fractional integration and dual integral equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):685–693, ???? 1962. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rudin:1962:SCW

[182] Walter Rudin. A strong converse of the Wiener–Levy theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 694–701, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdos:1962:CCG

[183] P. Erdős and C. A. Rogers. The construction of certain graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 702–707, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1962:CS

[184] W. T. Tutte. A census of slicings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14 (??):708–722, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sledd:1962:GPL

[185] W. T. Sledd. The Gibbs phenomenon and Lebesgue constants for regular Sonnenschein matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??): 723–728, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dowidar:1963:DSS

[186] A. F. Dowidar and G. M. Peterson. The distribution of sequences and summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):1–10, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ray-Chaudhuri:1963:AMC

[187] D. K. Ray-Chaudhuri. An algorithm for a minimum cover of an abstract complex. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):11–24, ????

1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1963:NPC

[188] Marion B. Smith, Jr. A new proof of certain metrization theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 25–32, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nordhaus:1963:TOG

[189] E. A. Nordhaus and B. M. Stewart. Triangles in an ordinary graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 33–41, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butson:1963:RAG

[190] A. T. Butson. Relations among generalized Hadamard matrices, relative difference sets, and maximal length linear recurring sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):42–48, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bush:1963:ETM

[191] George C. Bush. The embedding theorems of Malcev and Lambek. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 49–58, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leicht:1963:CS

[192] Johann B. Leicht. On commutative squares. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):59–79, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herstein:1963:SSD

[193] I. N. Herstein and W. R. Scott. Subnormal subgroups of division rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 80–83, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dedecker:1963:CNA

[194] Paul Dedecker. Sur la cohomologie non abélienne. II. (French) [On nonabelian cohomology. II]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):84–93, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlitz:1963:DCD

[195] L. Carlitz and John Riordan. The divided central differences of zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 94–100, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Singal:1963:UCS

[196] M. K. Singal. Union curves in a subspace V_{n} of a Riemannian V_{m}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 101–105, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weinstein:1963:NDE

[197] Joseph M. Weinstein. On the number of disjoint edges in a graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 106–111, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Church:1963:CSF

[198] P. T. Church. Cluster sets of functions on an n-ball. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):112–120, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Newman:1963:MDS

[199] Morris Newman. Multipliers of difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):121–124, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Browder:1963:EWT

[200] Felix E. Browder. On the “edge of the wedge” theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):125–131, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levy:1963:TFD

[201] Lawrence Levy. Torsion-free and divisible modules over non-integral-domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):132–151, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stamey:1963:UTC

[202] William L. Stamey and J. M. Marr. Unions of two convex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 152–156, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mitchell:1963:PHF

[203] Josephine Mitchell. Properties of harmonic functions of three real variables given by Bergman–Whittaker operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):157–168, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mills:1963:CPV

[204] W. H. Mills. Characters with preassigned values. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):169–171, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehmer:1963:PCP

[205] D. H. Lehmer, Emma Lehmer, and W. H. Mills. Pairs of consecutive power residues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):172–177, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jiggs:1963:RRC

[206] B. H. Jiggs. Recent results in commafree codes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):178–187, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haber:1963:MTR

[207] Robert M. Haber. Minimal term rank of a class of (0, 1)-matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 188–192, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wolf:1963:SFG

[208] Joseph A. Wolf. Space forms of Grassmann manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):193–205, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Motzkin:1963:RPL

[209] T. S. Motzkin. Rational points on linear subspaces. representation of an integer as a sum of squares with accessory conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):206–213, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mahowald:1963:SF

[210] Mark Mahowald. On singular fiberings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):214–219, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thompson:1963:NNM

[211] R. C. Thompson. A note on normal matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):220–225, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1963:MFL

[212] H. W. Ellis and Hidegorô Nakano. Monotone functions on linear lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):226–236, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harary:1963:EBG

[213] Frank Harary and Geert Prins. Enumeration of bicolourable graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 237–248, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1963:CPM

[214] W. T. Tutte. A census of planar maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):249–271, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Foulkes:1963:RGR

[215] H. O. Foulkes. On Redfield’s group reduction functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):272–284, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taft:1963:CSA

[216] Earl J. Taft. Cayley symmetries in associative algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):285–290, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chalk:1963:NSC

[217] J. H. H. Chalk. The number of solutions of congruences in incomplete residue systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):291–296, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1963:PRI

[218] R. E. Johnson. Principal right ideal rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):297–301, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wonenburger:1963:AGR

[219] María J. Wonenburger. The automorphisms of the group of rotations and its projective group corresponding to quadratic forms of any index. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 302–303, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlitz:1963:NCG

[220] L. Carlitz. The number of coloured graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):304–312, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wigner:1963:WPM

[221] Eugene P. Wigner. On weakly positive matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):313–317, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bottema:1963:CRT

[222] O. Bottema. On certain remarkable triangles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):318–322, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pearcy:1963:UEM

[223] Carl Pearcy. On unitary equivalence of matrices over the ring of continuous complex-valued functions on a Stonian space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):323–331, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wyman:1963:ABH

[224] Max Wyman. The asymptotic behaviour of the Hermite polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):332–349, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Federico:1963:NSL

[225] P. J. Federico. Note on some low-order perfect squares. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):350–362, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lambek:1963:URQ

[226] Joachim Lambek. On Utumi’s ring of quotients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):363–370, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fulkerson:1963:WSS

[227] D. R. Fulkerson and H. J. Ryser. Width sequences for special classes of (0, 1)-matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):371–396, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

LHeureux:1963:CCB

[228] James E. L’Heureux. Characterization of certain binary relations on connected ordered spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):397–411, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chalk:1963:IBQ

[229] J. H. H. Chalk. Integral bases for quadratic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):412–421, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clements:1963:ESF

[230] G. F. Clements. Entropies of sets of functions of bounded variation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 422–432, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Henstock:1963:TTI

[231] Ralph Henstock. Tauberian theorems for integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):433–439, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wright:1963:BA

[232] Fred B. Wright. Boolean averages. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):440–455, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cullen:1963:IFM

[233] C. G. Cullen. Intrinsic functions on matrices of real quaternions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 456–466, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yen:1963:ISR

[234] Ti Yen. Isomorphisms of simple rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):467–470, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cargo:1963:RAL

[235] G. T. Cargo. Radial and angular limits of meromorphic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 471–474, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Golomb:1963:SRF

[236] Solomon W. Golomb. On the sum of the reciprocals of the Fermat numbers and related irrationalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 475–478, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nash-Williams:1963:DGT

[237] C. St. J. A. Nash-Williams. Decomposition of graphs into two-way infinite paths. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):479–485, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:1963:CF

[238] Yung-Ming Chen. On conjugate functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):486–494, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alexiewicz:1963:SAS

[239] A. Alexiewicz and M. G. Arsove. Simultaneous automorphisms in spaces of analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):495–502, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miracle:1963:SRM

[240] Chester L. Miracle. Some regular [F, d_{n}] matrices with complex elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):503–525, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1963:ENS

[241] William G. Brown. Enumeration of non-separable planar maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 526–545, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lane:1963:PD

[242] N. D. Lane. Parabolic differentiation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):546–562, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deBranges:1963:SAS

[243] Louis de Branges. Some applications of spaces of entire functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 563–583, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Evans:1963:DPT

[244] Trevor Evans. A decision problem for transformations of trees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 584–590, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Finney:1963:PLS

[245] Ross Finney. Point-like, simplicial mappings of a 3-sphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 591–604, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banaschewski:1963:CRF

[246] B. Banaschewski. On the character rings of finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 605–612, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Snow:1963:MVV

[247] D. O. Snow. On measurability for vector-valued functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 613–621, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohn:1963:GO

[248] James A. Cohn and Donald Livingstone. On groups of order p^{3}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 622–624, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jones:1963:IRD

[249] Alfredo Jones. Integral representations of the direct product of groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 625–630, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Love:1963:DIE

[250] E. R. Love. Dual integral equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):631–640, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davies:1963:BFA

[251] I. J. Davies. A bipartitional function arising in Hall’s algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 641–649, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Connell:1963:GR

[252] Ian G. Connell. On the group ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):650–685, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shrikhande:1963:NES

[253] S. S. Shrikhande, D. Raghavarao, and S. K. Tharthare. Non-existence of some unsymmetrical partially balanced incomplete block designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??): 686–701, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [1012].

Hanani:1963:STC

[254] Haim Hanani. On some tactical configurations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):702–722, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Murdoch:1963:SMR

[255] D. C. Murdoch. Subrings of the maximal ring of quotients associated with closure operations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):723–743, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grunbaum:1963:NHS

[256] B. Grünbaum and T. S. Motzkin. The number of hexagons and the simplicity of geodesics on certain polyhedra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):744–751, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Appling:1963:IFN

[257] William D. L. Appling. Interval functions and non-decreasing functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):752–754, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCulloh:1963:IBK

[258] Leon R. McCulloh. Integral bases in Kummer extensions of Dedekind fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15 (??):755–765, ???? 1963. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chehata:1963:ETA

[259] C. G. Chehata. Embedding theorems for abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 15(??):766–770, ???? 1963. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ceder:1964:OSL

[260] Jack Ceder. On outwardly simple line families. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):1–11, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gale:1964:NFC

[261] David Gale. On the number of faces of a convex polytope. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):12–17, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mallory:1964:CS

[262] Donald J. Mallory and Maurice Sion. On covering systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):18–27, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yood:1964:ITR

[263] Bertram Yood. Ideals in topological rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):28–45, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pierce:1964:CAM

[264] William A. Pierce. Collineations of affine Moulton planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):46–62, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Killgrove:1964:CQP

[265] R. B. Killgrove. Completions of quadrangles in projective planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 63–76, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ginzburg:1964:NCL

[266] A. Ginzburg. A note on Cayley loops. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16 (??):77–81, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mesner:1964:TCM

[267] Dale M. Mesner. Traces of a class of (0, 1)-matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):82–93, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cantor:1964:DSC

[268] David G. Cantor. Determining a set from the cardinalities of its intersections with other sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):94–97, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hirschman:1964:VDT

[269] I. I. Hirschman, Jr. Variationdiminishing transformations and general orthogonal polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 98–107, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1964:MG

[270] W. T. Tutte. From matrices to graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16 (??):108–127, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wright:1964:CCG

[271] E. M. Wright. Counting coloured graphs. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):128–135, ????

1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dean:1964:FLG

[272] R. A. Dean. Free lattices generated by partially ordered sets and preserving bounds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):136–148, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Olubummo:1964:CBA

[273] A. Olubummo. Complemented Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):149–150, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gordon:1964:SDF

[274] B. Gordon and K. Rogers. Sums of the divisor function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):151–158, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Keller:1964:MMR

[275] Roy F. Keller. Motions of matrix rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16 (??):159–165, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kostant:1964:HEA

[276] B. Kostant and A. Novikoff. A homomorphism in exterior algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 166–168, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lane:1964:CD

[277] N. D. Lane and K. D. Singh. Conical differentiation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):169–190, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1964:MR

[278] G. de B. Robinson. Modular representations of sn. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):191–203, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

James:1964:WCS

[279] Robert C. James. Weak compactness and separation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):204–206, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bressler:1964:CTA

[280] D. W. Bressler and M. Sion. The current theory of analytic sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 207–230, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cargo:1964:ABH

[281] G. T. Cargo. Almost-bounded holomorphic functions with prescribed ambiguous points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):231–240, ????

1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cheney:1964:BPS

[282] E. W. Cheney and A. Sharma. Bernstein power series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):241–252, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:1964:EDS

[283] Henry B. Cohen. The k-extremally disconnected spaces as projectives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 253–260, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Craig:1964:EFP

[284] Robert T. Craig. Extension of finite projective planes. I. uniform Hjelmslev planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):261–266, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dixon:1964:CRI

[285] John D. Dixon. Complete reducibility of infinite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):267–274, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edwards:1964:CBL

[286] R. E. Edwards. Convolutions as bilinear and linear operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):

275–285, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilbert:1964:RPS

[287] E. N. Gilbert. Randomly packed and solidly packed spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 286–298, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1964:GRB

[288] I. M. Isaacs and D. S. Passman. Groups with representations of bounded degree. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):299–309, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jordan:1964:PCP

[289] J. H. Jordan. Pairs of consecutive power residues or non-residues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 310–314, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Khan:1964:LRB

[290] Nisar A. Khan. Linear relations between higher matrix commutators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16 (??):315–320, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lane:1964:APO

[291] N. D. Lane. Arcs of parabolic order four. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):321–338, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meir:1964:TPC

[292] Amram Meir. On two problems concerning the generalized Lototsky transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):339–342, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muskat:1964:CSC

[293] Joseph B. Muskat. Criteria for solvability of certain congruences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 343–352, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Quintas:1964:ODC

[294] Louis V. Quintas. One-dimensional cell complexes with homeotopy group equal to zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):353–357, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1964:GHS

[295] P. G. Rooney. A generalization of the Hardy spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):358–369, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sagle:1964:ACA

[296] Arthur A. Sagle. On anti-commutative algebras with an invariant form. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 370–378, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sandler:1964:SFP

[297] Reuben Sandler. On subplanes of free planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):379–385, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Turyn:1964:MTD

[298] Richard J. Turyn. The multiplier theorem for difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 386–388, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

White:1964:LT

[299] G. K. White. Lattice tetrahedra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 389–396, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wigner:1964:PSN

[300] Eugene P. Wigner and Mutsuo M. Yanase. On the positive semidefinite nature of a certain matrix expression. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16 (??):397–406, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Traylor:1964:CMP

[301] D. R. Traylor. Concerning metrizability of pointwise paracompact Moore spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):407–411, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilbert:1964:TFP

[302] Richard C. Gilbert and Vernon A. Kramer. Trace formulas for powers of a Sturm–Liouville operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 412–422, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cowling:1964:CRS

[303] V. F. Cowling and C. L. Miracle. Corrections to and remarks on “Some results for the generalized Lototsky transform”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):423–428, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [156].

Mott:1964:ECR

[304] Joe Leonard Mott. Equivalent conditions for a ring to be a multiplication ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):429–434, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kohler:1964:FGA

[305] Joseph Kohler. Finite groups with all maximal subgroups of prime or prime square index. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):435–442, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Foulser:1964:FTC

[306] David A. Foulser. The flag-transitive collineation groups of the finite Desarguesian affine planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 443–472, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1964:CSS

[307] Marvin Marcus and Adil Yaqub. Compounds of skew-symmetric matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16 (??):473–478, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Laha:1964:DCF

[308] R. G. Laha. On the decomposition of a class of functions of bounded variation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16 (??):479–484, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moon:1964:TGA

[309] J. W. Moon. Tournaments with a given automorphism group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 485–489, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ree:1964:CCS

[310] Rimhak Ree. Construction of certain semi-simple groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):490–508, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mendelsohn:1964:ASW

[311] N. S. Mendelsohn. An algorithmic solution for a word problem in group theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):509–516, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Klee:1964:CAP

[312] Victor Klee. A combinatorial analogue of Poincaré’s duality theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 517–531, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stevens:1964:LHE

[313] Harlan Stevens. Linear homogeneous equations over finite rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 532–538, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilmore:1964:CCG

[314] P. C. Gilmore and A. J. Hoffman. A characterization of comparability graphs and of interval graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):

539–548, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vanstone:1964:CSG

[315] J. R. Vanstone. Connections satisfying a generalized Ricci lemma. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 549–560, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gottlieb:1964:HIP

[316] Daniel H. Gottlieb. Homotopy and isotopy properties of topological spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16 (??):561–571, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1964:ERN

[317] W. G. Brown and W. T. Tutte. On the enumeration of rooted non-separable planar maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):572–577, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fox:1964:SIE

[318] Charles Fox. The solution of an integral equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):578–586, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Puttaswamaiah:1964:IRA

[319] B. M. Puttaswamaiah and G. de B. Robinson. Induced representations and alternating groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):587–601, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Klee:1964:DPG

[320] Victor Klee. Diameters of polyhedral graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):602–614, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hanani:1964:CBI

[321] Haim Hanani. On covering of balanced incomplete block designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 615–625, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roberts:1964:RBD

[322] J. B. Roberts. Relations between the digits of numbers and equal sums of like powers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):626–636, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pierce:1964:CPM

[323] William A. Pierce. Collineations of projective Moulton planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 637–656, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leech:1964:SSP

[324] John Leech. Some sphere packings in higher space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):657–682, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schwerdtfeger:1964:PGO

[325] Hans Schwerdtfeger. Projective geometry in the one-dimensional affine group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16 (??):683–700, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Klee:1964:NVC

[326] Victor Klee. On the number of vertices of a convex polytope. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 701–720, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Forelli:1964:I

[327] Frank Forelli. The isometries of H^{p}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16 (??):721–728, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fulkerson:1964:MND

[328] D. R. Fulkerson. The maximum number of disjoint permutations contained in a matrix of zeros and ones. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 729–735, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shrikhande:1964:GHM

[329] S. S. Shrikhande. Generalized Hadamard matrices and orthogonal arrays of strength two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):736–740, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hill:1964:SSP

[330] J. D. Hill and W. T. Sledd. Summability-(Z, p) and sequences of periodic type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):741–754, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Raney:1964:FS

[331] George N. Raney. A formal solution of sum_{i=1}^{∞}A_{i}e^{B_{i}}X = X. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 755–762, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heath:1964:SPP

[332] R. W. Heath. Screenability, pointwise paracompactness, and metrization of Moore spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):763–770, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herstein:1964:NRS

[333] I. N. Herstein and L. Small. Nil rings satisfying certain chain conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16 (??):771–776, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zirakzadeh:1964:PTI

[334] A. Zirakzadeh. A property of a triangle inscribed in a convex curve. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 777–786, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnsen:1964:IMA

[335] E. C. Johnsen. The inverse multiplier for abelian group difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??): 787–796, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ree:1964:FGP

[336] Rimhak Ree. Sur une famille de groupes de permutations doublement transitifs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):797–820, ???? 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sabidussi:1964:IEG

[337] Gert Sabidussi. Infinite Euler graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16 (??):821–838, ???? 1964. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnsen:1965:MRC

[338] E. C. Johnsen. Matrix rational completions satisfying generalized incidence equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):1–12, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anvari:1965:DSS

[339] Morteza Anvari. Derivative of singular set-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):13–16, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bullen:1965:GPI

[340] P. S. Bullen. A general Perron integral. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):17–30, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taylor:1965:CTL

[341] John C. Taylor. A class of translation lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):31–39, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Foulis:1965:SCO

[342] D. J. Foulis. Semigroups coordinatizing orthomodular geometries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17 (??):40–51, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kimura:1965:SWA

[343] Naoki Kimura, Takayuki Tamura, and Rudolph Merkel. Semigroups in which all subsemigroups are left ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 52–62, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eggleston:1965:PTC

[344] H. G. Eggleston and A. Zirakzadeh. A property of two chords which divide a convex curve into four arcs of equal length. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):63–77, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osborn:1965:INA

[345] J. Marshall Osborn. Identities of nonassociative algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):78–92, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Toth:1965:TAF

[346] L. Fejes Tóth. On the total area of the faces of a four-dimensional polytope. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17 (??):93–99, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harrison:1965:CFA

[347] Michael A. Harrison. A census of finite automata. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):100–113, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ray-Chaudhuri:1965:SCF

[348] D. K. Ray-Chaudhuri. Some configurations in finite projective spaces and partially balanced incomplete block designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):114–123, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Unni:1965:GPD

[349] K. R. Unni. On generalized powers of the difference operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 124–129, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Outcalt:1965:ECA

[350] D. L. Outcalt. An extension of the class of alternative rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):130–141, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hayes:1965:FTC

[351] Charles A. Hayes, Jr. Fubini theorem for a certain type of integral. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 142–154, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Killgrove:1965:CQP

[352] Raymond B. Killgrove. Completions of quadrangles in projective planes. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17 (??):155–165, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fulkerson:1965:SPG

[353] D. R. Fulkerson, A. J. Hoffman, and M. H. McAndrew. Some properties of graphs with multiple edges. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 166–177, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barnes:1965:IDN

[354] Wilfred E. Barnes and William M. Cunnea. Ideal decompositions in Noetherian rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):178–184, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Castonguay:1965:DTA

[355] C. E. Castonguay and H. G. Helfenstein. A distortion theorem for analytic maps of annuli. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):185–198, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thierrin:1965:AMF

[356] G. Thierrin. Anneaux métaprimitifs. (French) [Metaprimitive rings]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 199–206, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McArthur:1965:UES

[357] Charles W. McArthur and James R. Retherford. Uniform and equicontinuous Schauder bases of subspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):

207–212, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

L:1965:VFI

[358] Hilt Arthur L. and Chuan chih Hsiung. Vector fields and infinitesimal transformations on almost-Hermitian manifolds with boundary. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):213–238, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Peleg:1965:KQG

[359] Bezalel Peleg. The kernel of mquota games. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):239–244, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carroll:1965:DCP

[360] R. W. Carroll and C. L. Wang. On the degenerate Cauchy problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 245–256, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Appling:1965:RUI

[361] William D. L. Appling. Refinementunbounded interval functions and absolute continuity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):257–260, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlitz:1965:DIP

[362] L. Carlitz. The distribution of irreducible polynomials in several indeterminates. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):261–266, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butts:1965:AMR

[363] H. S. Butts and R. C. Phillips. Almost multiplication rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):267–277, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Divinsky:1965:KRR

[364] N. Divinsky and A. Suliński. Kurosh radicals of rings with operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 278–280, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilmer:1965:CLI

[365] Robert W. Gilmer, Jr. The cancellation law for ideals in a commutative ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17 (??):281–287, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meir:1965:TEC

[366] A. Meir. Tauberian estimates concerning the regular Hausdorff and [J, f(x)] transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):288–301, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1965:EQD

[367] William G. Brown. Enumeration of quadrangular dissections of the disk. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17 (??):302–317, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dulmage:1965:SPN

[368] A. L. Dulmage and N. S. Mendelsohn. The structure of powers of nonnegative matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):318–330, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jonsson:1965:HPP

[369] W. Jónsson. (C, γ, μ)-homogeneity of projective planes and polarities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 331–334, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Netanyahu:1965:EPS

[370] E. Netanyahu. Extremal problems for Schlicht functions in the exterior of the unit circle. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):335–341, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kay:1965:CCP

[371] David C. Kay and Gary Chartrand. A characterization of certain Ptolemaic graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):342–346, ????

1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdos:1965:ICC

[372] P. Erdős and L. Pósa. On independent circuits contained in a graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 347–352, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorimer:1965:GCF

[373] P. J. Lorimer. T_{2}-groups and a characterization of the finite groups of Moebius transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 353–366, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Browder:1965:TBL

[374] Felix E. Browder. On a theorem of Beurling and Livingston. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 367–372, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mullin:1965:CRT

[375] R. C. Mullin. On counting rooted triangular maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):373–382, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edrei:1965:EMF

[376] Albert Edrei and Wolfgang H. J. Fuchs. Entire and meromorphic functions with asymptotically prescribed characteristic. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):383–395, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clunie:1965:IFH

[377] J. Clunie. On integral functions having prescribed asymptotic growth. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 396–404, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gallagher:1965:GLC

[378] P. X. Gallagher. The generation of the lower central series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):405–410, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Topping:1965:SHP

[379] David M. Topping. Some homological pathology in vector lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 411–428, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chehata:1965:CCP

[380] C. G. Chehata. On commutative continuation of partial endomorphisms of groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):429–433, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1965:LOR

[381] F. W. Anderson. Lattice-ordered rings of quotients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):434–448, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edmonds:1965:PTF

[382] Jack Edmonds. Paths, trees, and flowers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):449–467, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rothschild:1965:DGF

[383] Bruce Rothschild. The decomposition of graphs into a finite number of paths. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17 (??):468–479, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bercov:1965:DTC

[384] Ronald D. Bercov. The double transitivity of a class of permutation groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17 (??):480–493, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beineke:1965:GC

[385] Lowell W. Beineke and Frank Harary. The genus of the n-cube. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 494–496, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chakerian:1965:MSC

[386] G. D. Chakerian and S. K. Stein. On measures of symmetry of convex bodies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):497–504, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mendelsohn:1965:CAS

[387] N. S. Mendelsohn. Correction to: “An algorithmic solution for a word problem in group theory”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):505, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1965:MS

[388] Gaston Smith. On the (f, d_{n})-method of summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):506–526, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1965:MCH

[389] D. W. Robinson. On matrix commutators of higher order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):527–532, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Motzkin:1965:MGN

[390] T. S. Motzkin and E. G. Straus. Maxima for graphs and a new proof of a theorem of Turán. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):533–540, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McFarland:1965:MDS

[391] Robert McFarland and H. B. Mann. On multipliers of difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 541–542, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morris:1965:SRS

[392] A. O. Morris. The spin representation of the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 543–549, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sagle:1965:ACA

[393] Arthur A. Sagle. On anti-commutative algebras and analytic loops. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 550–558, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kakita:1965:HCO

[394] Takao Kakita. Hyperbolic convolution operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):559–582, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohn:1965:SGA

[395] James A. Cohn and Donald Livingstone. On the structure of group algebras. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):583–593, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1965:HRA

[396] T. Anderson, N. Divinsky, and A. Suliński. Hereditary radicals in associative and alternative rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 594–603, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ernest:1965:NGA

[397] John Ernest. A new group algebra for locally compact groups. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 604–615, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baker:1965:SDI

[398] A. Baker. On some Diophantine inequalities involving the exponential function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):616–626, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Caldwell:1965:SRB

[399] W. V. Caldwell. Some relationships between Bers and Beltrami systems and linear elliptic systems of partial differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):627–642, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hasumi:1965:ISC

[400] Morisuke Hasumi and T. P. Srinivasan. Invariant subspaces of continuous functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):643–651, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdos:1965:TQ

[401] Paul Erdős and A. Sharma. On Tchebycheff quadrature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 652–658, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kotzig:1965:CTP

[402] Anton Kotzig. Colouring of trivalent polyhedra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):659–664, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuyk:1965:CFI

[403] Willem Kuyk. The construction of fields with infinite cyclic automorphism group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):665–668, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:1965:NCF

[404] Henry B. Cohen. The k-normal completion of function lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 669–675, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilbert:1965:LSC

[405] R. P. Gilbert. On the location of singularities of a class of elliptic partial differential equations in four variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17 (??):676–686, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoffman:1965:LGF

[406] A. J. Hoffman and D. K. Ray-Chaudhuri. On the line graph of a finite affine plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):687–694, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1965:COH

[407] Arlen Brown, P. R. Halmos, and Carl Pearcy. Commutators of operators on Hilbert space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):695–708, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Halperin:1965:RRR

[408] Israel Halperin. Regular rank rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17 (??):709–719, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hajnal:1965:TSG

[409] A. Hajnal. A theorem on ksaturated graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):720–724, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bantegnie:1965:ICR

[410] Robert Bantegnie. Sur l’indice de certains réseaux de R^{4} permis pour un octaèdre. (French) [On the index of certain networks of R^{4} permitted for an octahedron]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):725–730, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramanujacharyulu:1965:TTE

[411] C. Ramanujacharyulu. Trees and treeequivalent graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):731–733, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kerr-Lawson:1965:FDH

[412] A. Kerr-Lawson. A filter description of the homomorphisms of H^{∞}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 734–757, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maranda:1965:FC

[413] J.-M. Maranda. Formal categories. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17 (??):758–801, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zaidman:1965:CPW

[414] S. Zaidman. Convexity properties for weak solutions of some differential equations in Hilbert spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):

802–807, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Murdoch:1965:NWD

[415] B. H. Murdoch. A note on welldistributed sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):808– 810, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brauer:1965:SRL

[416] Fred Brauer. Some refinements of Lyapunov’s second method. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 811–819, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Worrell:1965:CDT

[417] J. M. Worrell, Jr. and H. H. Wicke. Characterizations of developable topological spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):820–830, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Walkup:1965:MIM

[418] David W. Walkup. Minimal interchanges of (0, 1)-matrices and disjoint circuits in a graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):831–838, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edwards:1965:BII

[419] R. E. Edwards. Bipositive and isometric isomorphisms of some convolution algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):839–846, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Helfenstein:1965:TS

[420] H. G. Helfenstein. Topological H-surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):847–849, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beineke:1965:TCG

[421] Lowell W. Beineke and Frank Harary. The thickness of the complete graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17 (??):850–859, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gorenstein:1965:FGW

[422] Daniel Gorenstein. Finite groups in which Sylow 2-subgroups are abelian and centralizers of involutions are solvable. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):860–906, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

ONeill:1965:IKI

[423] Barrett O’Neill. Isotropic and Kähler immersions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):907–915, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ostrom:1965:CHP

[424] T. G. Ostrom. A characterization of the Hughes planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):916–922, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alvarez:1965:UGR

[425] Laurence R. Alvarez. Undirected graphs realizable as graphs of modular lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):923–932, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bernstein:1965:PJP

[426] Leon Bernstein. A periodic Jacobi– Perron algorithm. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):933–945, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cunningham:1965:KC

[427] F. Cunningham, Jr. and I. J. Schoenberg. On the Kakeya constant. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 946–956, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fulkerson:1965:URR

[428] D. R. Fulkerson. Upsets in round robin tournaments. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):957–969, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gioia:1965:PAF

[429] A. A. Gioia. The K-product of arithmetic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):970–976, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rigby:1965:ASF

[430] J. F. Rigby. Affine subplanes of finite projective planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):977–1009, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gordon:1965:TTM

[431] B. Gordon and M. M. Robertson. Two theorems on mosaics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):1010– 1014, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akasaki:1965:IMS

[432] Takeo Akasaki. Immersions of metric spaces into Euclidean spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 1015–1018, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Webb:1965:SRN

[433] W. A. Webb. Sums of rational numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):1019–1024, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Albert:1965:NFR

[434] A. A. Albert. A normal form for Riemann matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):1025–1029, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coddington:1965:FNO

[435] Earl A. Coddington. Formally normal operators having no normal extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17 (??):1030–1040, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feller:1965:PM

[436] E. H. Feller and E. W. Swokowski. Prime modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):1041–1052, ???? 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Swanson:1965:DPB

[437] C. A. Swanson. Domain perturbations of the biharmonic operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??): 1053–1063, ???? 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Majindar:1966:IMIa

[438] Kulendra N. Majindar. On integer matrices and incidence matrices of certain combinatorial configurations. I. square matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1–5, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Majindar:1966:IMIb

[439] Kulendra N. Majindar. On integer matrices and incidence matrices of certain combinatorial configurations. II. rectangular matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):6–8, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Majindar:1966:IMIc

[440] Kulendra N. Majindar. On integer matrices and incidence matrices of certain combinatorial configurations. III. rectangular matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):9–17, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lewin:1966:SFG

[441] Jacques Lewin. Some finitely generated analogues of a group of A. H. Clifford. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):18–26, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cross:1966:FFS

[442] George Cross. A Fourier formula for series summable (C, 1). Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):27–32, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mullin:1966:ANT

[443] R. C. Mullin. On the average number of trees in certain maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):33–41, ????

1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cantor:1966:DSC

[444] David G. Cantor and W. H. Mills. Determination of a subset from certain combinatorial properties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 42–48, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Henstock:1966:MVI

[445] Ralph Henstock. Majorants in variational integration. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):49–74, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tang:1966:ETG

[446] C. Y. Tang. An existence theorem for generalized direct products with amalgamated subgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):75–82, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gaifman:1966:RCL

[447] Haim Gaifman. Remarks on complementation in the lattice of all topologies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):83–88, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

King:1966:LTB

[448] J. P. King. The Lototsky transform and Bernstein polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 89–91, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wall:1966:AIS

[449] C. T. C. Wall. Arithmetic invariants of subdivision of complexes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 92–96, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Segal:1966:ISM

[450] S. L. Segal. On Ingham’s summation method. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):97–105, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdos:1966:RGS

[451] Paul Erdős, A. W. Goodman, and Lajos Pósa. The representation of a graph by set intersections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):106–112, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Horadam:1966:GAC

[452] A. F. Horadam. Groups associated with certain loci in [5]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 113–119, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bryant:1966:PLI

[453] B. F. Bryant and Hans Schneider. Principal loop-isotopes of quasigroups.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):120–125, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brualdi:1966:TRD

[454] Richard A. Brualdi. Term rank of the direct product of matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 126–138, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cullen:1966:CFA

[455] C. G. Cullen and C. A. Hall. Classes of functions on algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 139–146, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Scott:1966:CPE

[456] P. R. Scott. The construction of perfect and extreme forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):147–158, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sherman:1966:WTU

[457] Thomas Sherman. A weight theory for unitary representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 159–168, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1966:CPR

[458] Norman W. Johnson. Convex polyhedra with regular faces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 169–200, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Copeland:1966:IDC

[459] Arthur H. Copeland, Jr. Isotopy of 2dimensional cones. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):201–210, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hemminger:1966:GDG

[460] Robert L. Hemminger. On the group of a directed graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):211–220, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meir:1966:FNL

[461] A. Meir. A further note on lototskytype transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):221– 224, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DiPaola:1966:RCB

[462] Jane W. Di Paola. On a restricted class of block design games. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 225–236, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baum:1966:NCP

[463] L. E. Baum and J. A. Eagon. The number of circular patterns compatible with a pseudo-symmetric connected graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):237–239, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hasumi:1966:IST

[464] Morisuke Hasumi. Invariant subspace theorems for finite Riemann surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):240–255, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lappan:1966:NFN

[465] P. Lappan and D. C. Rung. Normal functions and non-tangential boundary arcs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):256–264, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hu:1966:DEQ

[466] Tah kai Hu. Distributive extensions and quasi-framal algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 265–281, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sagle:1966:SAG

[467] Arthur A. Sagle. Simple algebras that generalize the Jordan algebra M_{3}^{8}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):282–290, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wunderlich:1966:SGS

[468] M. C. Wunderlich. Sieve-generated sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):291–299, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herstein:1966:NRS

[469] I. N. Herstein and Lance Small. Nil rings satisfying certain chain conditions: An addendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):300– 302, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sinkhorn:1966:RBA

[470] Richard Sinkhorn. A relationship between arbitrary positive matrices and stochastic matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):303–306, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steger:1966:EFR

[471] Arthur Steger. Elementary factorization in π-regular rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 307–313, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hinton:1966:SVI

[472] D. B. Hinton. A Stieltjes–Volterra integral equation theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):314–331, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Raney:1966:GFS

[473] George N. Raney. Generalization of the Fibonacci sequence to n dimensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):332–349, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hurd:1966:BRS

[474] A. E. Hurd. Boundary regularity in the Sobolev imbedding theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 350–356, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benedetto:1966:LTG

[475] John Benedetto. The Laplace transform of generalized functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 357–374, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edelstein:1966:BFC

[476] M. Edelstein and L. M. Kelly. Bisecants of finite collections of sets in linear spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):375–380, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ginsburg:1966:CMM

[477] Seymour Ginsburg and Gene F. Rose. A characterization of machine mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):381–388, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rider:1966:GSG

[478] Daniel Rider. Gap series on groups and spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):389–398, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Voichick:1966:ISR

[479] Michael Voichick. Invariant subspaces on Riemann surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):399– 403, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Utumi:1966:CSI

[480] Yuzo Utumi. On the continuity and self-injectivity of a complete regular ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):404–412, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stamm:1966:SHC

[481] Emil Stamm. Simplicial and homotopical cohomology of polyhedra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 413–423, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DeMarr:1966:MCT

[482] Ralph DeMarr. A martingale convergence theorem in vector lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 424–432, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clark:1966:WSS

[483] W. Edwin Clark. Weakly semi-simple finite-dimensional algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 433–442, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barros-Neto:1966:SGK

[484] José Barros-Neto. On the smoothness of general kernels. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):443–448, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoehnke:1966:SS

[485] Hans-Jürgen Hoehnke. Structure of semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):449–491, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bacon:1966:EFB

[486] Philip Bacon. Equivalent formulations of the Borsuk–Ulam theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 492–502, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ditzian:1966:ITL

[487] Z. Ditzian and A. Jakimovski. Inversion theorems for the Laplace– Stieltjes transform. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):503–516, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McMillan:1966:LPE

[488] D. R. McMillan, Jr. Local properties of the embedding of a graph in a three-manifold. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):517–528, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grunbaum:1966:CFC

[489] Branko Grünbaum. Continuous families of curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):529–537, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hirschman:1966:SPC

[490] I. I. Hirschman, Jr. Szegő polynomials on a compact group with ordered dual. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):538–560, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Porter:1966:PBE

[491] A. Duane Porter. Pairs of bilinear equations in a finite field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):561–565, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barnes:1966:MAA

[492] Bruce A. Barnes. Modular annihilator algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):566–578, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glauberman:1966:ECG

[493] George Glauberman, Eugene F. Krause, and Ruth Rebekka Struik. Engel congruences in groups of primepower exponent. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):579–588, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wilson:1966:CLI

[494] Eric L. Wilson. A class of loops with the isotopy-isomorphy property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 589–592, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DeMar:1966:UCD

[495] Richard F. DeMar. Uniqueness classes for difference functionals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 593–607, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brualdi:1966:IPM

[496] R. A. Brualdi and M. Newman. Inequalities for the permanental minors of non-negative matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):

608–615, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Magill:1966:CC

[497] Kenneth D. Magill, Jr. Countable compactifications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):616–620, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Danicic:1966:IPL

[498] I. Danicic. On the integral part of a linear form with prime variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 621–628, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tondeur:1966:ISB

[499] Philippe Tondeur. Invariant subbundles of the tangent bundle of a homogeneous space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):629–634, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berg:1966:NCF

[500] I. D. Berg. A note on convergence fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):635–638, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shatz:1966:NDT

[501] Stephen S. Shatz. Note on the descendent theorem of Slepian, Moore, and Prange. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):639–642, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Folkman:1966:RIS

[502] Jon Folkman. On the representation of integers as sums of distinct terms from a fixed sequence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):643–655, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heyde:1966:SRS

[503] C. C. Heyde. Some results on small-deviation probability convergence rates for sums of independent random variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):656–665, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ostrom:1966:RNN

[504] T. G. Ostrom. Replaceable nets, net collineations, and net extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):666–672, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Peleg:1966:KGS

[505] B. Peleg. The kernel of the general-sum four-person game. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):673–677, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Quintas:1966:EST

[506] Louis V. Quintas and Fred Supnick. Extrema in space-time. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 678–691, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nash-Williams:1966:ELI

[507] C. St. J. A. Nash-Williams. Euler lines in infinite directed graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 692–714, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fearnley:1966:ETP

[508] L. Fearnley. Embeddings of topological products of circularly chainable continua. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):715–723, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gross:1966:PCE

[509] Fred Gross. On the periodicity of compositions of entire functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 724–730, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barnette:1966:TPG

[510] David Barnette. Trees in polyhedral graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):731–736, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1966:MCO

[511] Arlen Brown and Carl Pearcy. Multiplicative commutators of operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):737–749, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rolfsen:1966:AMP

[512] Dale Rolfsen. Alternative metrization proofs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):750–757, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wilf:1966:PDS

[513] Herbert S. Wilf. On the permanent of a doubly stochastic matrix. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 758–761, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deLamadrid:1966:MTI

[514] Jesús Gil de Lamadrid. Measures and tensors. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):762–793, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haimo:1966:SEG

[515] Deborah Tepper Haimo. Series expansions of generalized temperature functions in N dimensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 794–802, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harary:1966:RTM

[516] Frank Harary and Ed Palmer. The reconstruction of a tree from its maximal subtrees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):803–810, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robkin:1966:PLA

[517] Eugene Robkin and F. A. Valentine. Parallel lines associated with two sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):811–816, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duke:1966:GRN

[518] Richard A. Duke. The genus, regional number, and Betti number of a graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):817–822, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feller:1966:SPM

[519] E. H. Feller and E. W. Swokowski. Semi-prime modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):823– 831, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rinehart:1966:MF

[520] R. F. Rinehart. P and D in P^{−1}XP = dg(λ_{1}, ⋯, λ_{n}) = D as matrix functions of X. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):832–837, ????

1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Melzak:1966:IPD

[521] Z. A. Melzak. Infinite packings of disks. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):838–852, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harary:1966:ELR

[522] Frank Harary and Ed Palmer. Enumeration of locally restricted digraphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):853–860, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ton:1966:SPN

[523] Bui An Ton. Singular perturbations of nonlinear elliptic and parabolic variational boundary-value problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 861–872, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Whittaker:1966:MCP

[524] James V. Whittaker. A mountainclimbing problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):873–882, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McKinney:1966:UTC

[525] Richard L. McKinney. On unions of two convex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):883–886, ????

1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nakai:1966:CPF

[526] Mitsuru Nakai and Leo Sario. Construction of principal functions by orthogonal projection. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):887–896, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fillmore:1966:PS

[527] Peter A. Fillmore. On products of symmetries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):897–900, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Helzer:1966:DI

[528] Garry Helzer. On divisibility and injectivity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):901–919, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Trojan:1966:IEI

[529] Allan Trojan. The integral extension of isometries of quadratic forms over local fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):920–942, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jensen:1966:RFP

[530] Chr. U. Jensen. A remark on flat and projective modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):943–949, ????

1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCarthy:1966:NPI

[531] P. J. McCarthy. Note on primary ideal decompositions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):950–952, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Courter:1966:MCR

[532] R. C. Courter. The maximal corational extension by a module. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 953–962, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Freese:1966:CBM

[533] Raymond W. Freese and Edward Z. Andalafte. A characterization of 2betweenness in 2-metric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 963–968, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Everett:1966:BCN

[534] C. J. Everett. On the basis and chromatic number of a graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 969–973, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kestelman:1966:SCS

[535] H. Kestelman. Self-centered sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):974–980, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shapiro:1966:EP

[536] H. L. Shapiro. Extensions of pseudometrics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):981–998, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stewart:1966:SFB

[537] B. M. Stewart and W. A. Webb. Sums of fractions with bounded numerators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):999–1003, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alo:1966:TLP

[538] Richard A. Alò and Orrin Frink. Topologies of lattice products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 1004–1014, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hayes:1966:CHT

[539] C. A. Hayes, Jr. A condition of halo type for the differentiation of classes of integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1015–1023, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ohm:1966:PIP

[540] Jack Ohm. Primary ideals in Prüfer domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1024–1030, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stewart:1966:MG

[541] B. M. Stewart. Magic graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 1031–1059, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Foulkes:1966:RRC

[542] H. O. Foulkes. On Redfield’s rangecorrespondences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1060–1071, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Linden:1966:ZTB

[543] C. N. Linden and H. Somadasa. Zero tracts of Blaschke products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 1072–1078, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clark:1966:HSP

[544] Colin Clark. The Hilbert–Schmidt property for embedding maps between Sobolev spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1079–1084, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1966:PIC

[545] Marvin Marcus and Henryk Minc. A permanental inequality—the case of equality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1085–1090, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benson:1966:MRG

[546] Clark T. Benson. Minimal regular graphs of girths eight and twelve. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):1091–1094, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wong:1966:GCC

[547] James S. W. Wong. Generalizations of the converse of the contraction mapping principle. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1095–1104, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adams:1966:QLE

[548] R. A. Adams. A quasi-linear elliptic boundary value problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 1105–1112, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:1966:SMV

[549] Q. I. Rahman. The second mean values of entire functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1113–1120, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eisen:1966:PAS

[550] M. Eisen. Piecewise analytic solutions of mixed boundary value problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):1121–1147, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wu:1966:BIB

[551] Ling erl E. T. Wu. Bicontinuous isomorphisms between two closed left ideals of a compact dual ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 1148–1151, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1966:CFT

[552] Arlen Brown and Carl Pearcy. Commutators in factors of type III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 1152–1160, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bose:1966:HVF

[553] R. C. Bose and I. M. Chakravarti. Hermitian varieties in a finite projective space PG(N, q^{2}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1161–1182, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butts:1966:PIP

[554] H. S. Butts and Robert W. Gilmer, Jr. Primary ideals and prime power ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1183–1195, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Motzkin:1966:NIA

[555] T. S. Motzkin and A. Sharma. Nextto-interpolatory approximation on sets with multiplicities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1196–1211, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Janowitz:1966:SAL

[556] M. F. Janowitz. A semigroup approach to lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1212–1223, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zakon:1966:EMU

[557] Elias Zakon. On the “essential metrization” of uniform spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 1224–1236, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hedrlin:1966:RUG

[558] Z. Hedrlín and A. Pultr. On rigid undirected graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1237–1242, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1966:HTA

[559] I. M. Isaacs and D. S. Passman. Half-transitive automorphism groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):1243–1250, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Powell:1966:ST

[560] Robert E. Powell. The L(r, t) summability transform. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1251–1260, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jakimovski:1966:SHN

[561] Amnon Jakimovski. Summability of the Heine and Neumann series of Legendre polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1261–1263, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Friedman:1966:AFS

[562] N. Friedman and M. Katz. Additive functionals of L_{p} spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 1264–1271, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mitchell:1966:BSS

[563] Josephine Mitchell. Bounds for solutions of a system of linear partial differential equations on domains with Bergman–Silov boundary. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??): 1272–1280, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ruckle:1966:CSS

[564] William Ruckle. On the construction of sequence spaces that have Schauder bases. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1281–1293, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shephard:1966:SPC

[565] G. C. Shephard. The Steiner point of a convex polytope. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1294–1300, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1966:CM

[566] W. T. Tutte. Connectivity in matroids. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18 (??):1301–1324, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cooper:1966:LTD

[567] J. L. B. Cooper. Laplace transformations of distributions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1325– 1332, ???? 1966. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dawkins:1966:ICC

[568] Brian P. Dawkins and Israel Halperin. The isomorphism of certain continuous rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):1333–1344, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barback:1967:DSI

[569] Joseph Barback. Double series of isols. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):1–15, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sharma:1967:IPR

[570] A. Sharma. Interpolation by polynomials in z and z^{−1} in the roots of unity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):16–23, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aubert:1967:TAG

[571] K. E. Aubert. Un théorème sur l’arithmétique des groupes ordonnés. (French) [A theorem on the arithmetic of ordered groups]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):24–27, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Darsow:1967:SM

[572] William F. Darsow. Signal metrics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):28–39, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tewari:1967:CCA

[573] Krishna Tewari. Complexes over a complete algebra of quotients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 40–57, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Finlayson:1967:NAW

[574] Henry C. Finlayson. New approximations for Wiener integrals, with error estimates. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):58–105, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Labute:1967:CDG

[575] John P. Labute. Classification of Demushkin groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):106–132, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gulick:1967:GAMa

[576] S. L. Gulick, T. S. Liu, and A. C. M. van Rooij. Group algebra modules. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):133–150, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gulick:1967:GAMb

[577] S. L. Gulick, T. S. Liu, and A. C. M. van Rooij. Group algebra modules. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):151–173, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mullin:1967:ETR

[578] R. C. Mullin. On the enumeration of tree-rooted maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):174–183, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Howard:1967:OCM

[579] H. C. Howard. Oscillation criteria for matrix differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 184–199, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [983].

Rockafellar:1967:CLT

[580] R. T. Rockafellar. Conjugates and Legendre transforms of convex functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):200–205, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McEliece:1967:GDS

[581] Robert J. McEliece. A generalization of difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):206–211, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barbeau:1967:SPS

[582] Edward J. Barbeau. Subordinate and pseudo-subordinate semi-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):212–224, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1967:DIR

[583] H. W. Ellis and R. L. Jeffery. Derivatives and integrals with respect to a base function of generalized bounded variation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):225–241, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ingram:1967:CNP

[584] W. T. Ingram. Concerning non-planar circle-like continua. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):242–250, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leech:1967:NSP

[585] John Leech. Notes on sphere packings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):251–267, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McQuillan:1967:NWP

[586] Donald L. McQuillan. A note on Weierstrass points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):268–272, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mesner:1967:SDL

[587] Dale M. Mesner. Sets of disjoint lines in PG(3, q). Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):273–280, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Botta:1967:LTM

[588] P. Botta. Linear transformations on matrices: The invariance of a class of general matrix functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 281–290, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McLaughlin:1967:SDR

[589] T. G. McLaughlin. Splitting and decomposition by regressive sets. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):291–311, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Forelli:1967:EP

[590] Frank Forelli. Extreme points in H^{1}(R). Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):312–320, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCord:1967:ELC

[591] Michael C. McCord. Embedding Plike compacta in manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 321–332, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barrar:1967:PFP

[592] Richard B. Barrar. Proof of the fixed point theorems of Poincaré and Birkhoff. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):333–343, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pearl:1967:DTM

[593] Martin H. Pearl. A decomposition theorem for matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):344–349, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Larson:1967:CHA

[594] Richard G. Larson. Cocommutative Hopf algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):350–360, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Akasaki:1967:KHM

[595] T. Akasaki. On a kind of homotopy manifold. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):361–365, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Peleg:1967:EPO

[596] Bezalel Peleg. Equilibrium points for open acyclic relations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 366–369, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fujiwara:1967:PSC

[597] Tsuyoshi Fujiwara. Permutability of semilattice congruences of lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):370–375, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:1967:AFP

[598] G. E. Martin. On arcs in a finite projective plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):376–393, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Graham:1967:PET

[599] R. L. Graham. On partitions of an equilateral triangle. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):394–409, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wehrfritz:1967:NG

[600] B. A. F. Wehrfritz. A note on d-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):410–412, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heppes:1967:NSW

[601] Aladár Heppes. On the number of spheres which can hide a given sphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):413–418, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Warne:1967:EBS

[602] R. J. Warne. Extensions of I-bisimple semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):419–426, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See errata [741].

Stewart:1967:SCG

[603] B. M. Stewart. Supermagic complete graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):427–438, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yano:1967:NHR

[604] Kentaro Yano. Notes on hypersurfaces in a Riemannian manifold. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 439–446, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dickson:1967:NHR

[605] S. E. Dickson. A note on hypernilpotent radical properties for associative rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):447–448, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Libera:1967:USF

[606] Richard J. Libera. Univalent αspiral functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):449–456, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bullen:1967:GPI

[607] P. S. Bullen. A general Perron integral. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):457–473, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borges:1967:ET

[608] Carlos J. R. Borges. On extensions of topologies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):474–487, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Luxemburg:1967:ECO

[609] W. A. J. Luxemburg and J. J. Masterson. An extension of the concept of the order dual of a Riesz space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 488–498, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putz:1967:TFB

[610] H. Putz. Triangulation of fibre bundles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):499–513, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jordan:1967:CCR

[611] James H. Jordan. Covering classes of residues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):514–519, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hobby:1967:NRG

[612] C. R. Hobby. Nearly regular p-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):520–522, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bollobas:1967:EPP

[613] Béla Bollobás. An extremal problem for polygons inscribed in a convex curve. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):523–528, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heerema:1967:IIR

[614] N. Heerema. Inertial isomorphisms of v-rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):529–539, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brickell:1967:DMA

[615] F. Brickell. Differentiable manifolds with an area measure. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):540– 549, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Csorgo:1967:NPS

[616] Miklós Csörgö. A new proof of some results of Rényi and the asymptotic distribution of the range of his Kolmogorov–Smirnov type random variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):550–558, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Buschman:1967:SGS

[617] R. G. Buschman and M. C. Wunderlich. Sieve-generated sequences with translated intervals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):559–570, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brauer:1967:GFS

[618] Fred Brauer. Green’s functions for singular ordinary differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):571–582, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Appel:1967:UGD

[619] K. I. Appel and E. T. Parker. On unsolvable groups of degree p = 4q + 1, p and q primes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):583–589, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hall:1967:IFS

[620] C. A. Hall. Intrinsic functions on semisimple algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):590–598, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boyd:1967:HTR

[621] D. W. Boyd. The Hilbert transform on rearrangement-invariant spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 599–616, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wong:1967:CFP

[622] James S. W. Wong. Common fixed points of commuting monotone mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):617–620, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meeker:1967:SGH

[623] L. D. Meeker. Some generalized Hardy spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):621–628, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mancuso:1967:ACN

[624] Vincent J. Mancuso. n-ANR’s for certain normal spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):629–635, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DeMarr:1967:POO

[625] Ralph DeMarr. On partially ordering operator algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):636–643, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1967:NET

[626] William G. Brown. On the nonexistence of a type of regular graphs of girth 5. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):644–648, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mack:1967:DCP

[627] John Mack. Directed covers and paracompact spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):649–654, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Williams:1967:PCR

[628] Kenneth S. Williams. Pairs of consecutive residues of polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):

655–666, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ogawa:1967:LBS

[629] Hajimu Ogawa. Lower bounds for solutions of parabolic differential inequalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):667–672, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartman:1967:HHM

[630] Philip Hartman. On homotopic harmonic maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):673–687, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dorroh:1967:SGM

[631] J. R. Dorroh. Semi-groups of maps in a locally compact space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 688–696, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zame:1967:DS

[632] Alan Zame. The distribution of sequences modulo1. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):697–709, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Griffin:1967:SRM

[633] Malcolm Griffin. Some results on vmultiplication rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):710–722, ????

1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Caradus:1967:MOIa

[634] S. R. Caradus. On meromorphic operators. I, II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):723–736, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Caradus:1967:MOIb

[635] S. R. Caradus. On meromorphic operators, II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):737–748, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sankoff:1967:PPM

[636] D. Sankoff and D. A. Dawson. A packing problem for measurable sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 749–756, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Luther:1967:UEP

[637] Norman Y. Luther. Unique extension and product measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 757–763, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nelson:1967:FSO

[638] Evelyn Nelson. Finiteness of semigroups of operators in universal algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):764–768, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jones:1967:DPN

[639] William B. Jones. Direct products of normed linear spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 769–773, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ginzburg:1967:RGG

[640] A. Ginzburg. Representation of groups by generalized normal multiplication tables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):774–791, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sheehan:1967:PT

[641] J. Sheehan. On Pólya’s theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 792–799, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harary:1967:IG

[642] Frank Harary and Michael D. Plummer. On indecomposable graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):800–809, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brickman:1967:ISL

[643] L. Brickman and P. A. Fillmore. The invariant subspace lattice of a linear transformation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):810–822, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Saxena:1967:ECD

[644] R. B. Saxena. Expansion of continuous differentiable functions in Fourier Legendre series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):823–827, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ruckle:1967:SCS

[645] William Ruckle. Symmetric coordinate spaces and symmetric bases. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 828–838, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boyd:1967:COL

[646] David W. Boyd. A class of operators on the Lorentz space M(phi). Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 839–841, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brockwell:1967:NAT

[647] P. J. Brockwell. On the number of associative triples in an algebra of n elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):842–850, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Klarner:1967:CGP

[648] David A. Klarner. Cell growth problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):851–863, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Oddson:1967:PLC

[649] J. K. Oddson. Phragmèn–Lindelöf and comparison theorems for ellipticparabolic differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 864–871, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Janko:1967:CFS

[650] Zvonimir Janko. A characterization of the finite simple group PSp_{4}, (3). Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):872–894, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bachmann:1967:HGI

[651] F. Bachmann. Der Höhensatz in der Geometrie involutorischer [Gruppenelemente]. (German) [The higher rate in involution geometry [group elements]]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):895–903, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baer:1967:SAG

[652] Reinhold Baer. Soluble Artinian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):904–923, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dembowski:1967:CGC

[653] Peter Dembowski. Collineation groups containing perspectivities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 924–937, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DuVal:1967:RSE

[654] Patrick Du Val. Rational surfaces with exceptional unodes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):938–951, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edge:1967:NLK

[655] W. L. Edge. A new look at the Kummer surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):952–967, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdos:1967:SAG

[656] P. Erdős. On some applications of graph theory to geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 968–971, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Firey:1967:GCB

[657] Wm. J. Firey. Generalized convex bodies of revolution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):972–996, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Forder:1967:OP

[658] H. G. Forder. On order in a plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):997–1000, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goethals:1967:OMZ

[659] J.-M. Goethals and J. J. Seidel. Orthogonal matrices with zero diagonal. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):1001–1010, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moser:1967:NVR

[660] W. O. J. Moser. The number of very reduced 4 × n Latin rectangles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 1011–1017, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Neumann:1967:SRS

[661] B. H. Neumann. Some remarks on semigroup presentations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 1018–1026, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum and addendum [740].

Rigby:1967:CCP

[662] J. F. Rigby. Collineations, correlations, polarities, and conics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):

1027–1041, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Scherk:1967:DAC

[663] Peter Scherk. On differentiable arcs and curves. VI. singular osculating spaces of curves of order n + 1 in projective n-space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1042–1061, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schoenberg:1967:CLC

[664] I. J. Schoenberg and S. C. Zaremba. On Cauchy’s lemma concerning convex polygons. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1062–1071, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schuster:1967:APC

[665] Seymour Schuster. Axes of points of certain linear systems of polarities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):1072–1077, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sherk:1967:FIS

[666] F. A. Sherk. Finite incidence structures with orthogonality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1078–1083, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Synge:1967:TIF

[667] J. L. Synge. Two isomorphs of the four-colour problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1084–1091, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Toth:1967:VPT

[668] L. Fejes Tóth and A. Heppes. A variant of the problem of the thirteen spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):1092–1100, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1967:AD

[669] W. T. Tutte. Antisymmetrical digraphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1101–1118, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alexandrov:1967:CC

[670] A. D. Alexandrov. A contribution to chronogeometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1119–1128, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bohm:1967:VNR

[671] J. Böhm. Zur Verallgemeinerung der Neperschen Regel in r-dimensionalen Riemannschen Räumen konstanter Krümmung. (German) [Generalization of the Napierian rule in rdimensional Riemannian spaces of constant curvature]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1129–1148, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bottema:1967:IDB

[672] O. Bottema. The inversive distance between two circles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1149–1152, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1967:DDD

[673] G. F. D. Duff. Differences, derivatives, and decreasing rearrangements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 1153–1178, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garner:1967:RSP

[674] Cyril W. L. Garner. Regular skew polyhedra in hyperbolic three-space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 1179–1186, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pedoe:1967:GTE

[675] Daniel Pedoe. On a geometrical theorem in exterior algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 1187–1191, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Macbeath:1967:CNE

[676] A. M. Macbeath. The classification of non-Euclidean plane crystallographic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1192–1205, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shephard:1967:PCS

[677] G. C. Shephard. Polytopes with centrally symmetric faces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1206– 1213, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shephard:1967:EPG

[678] G. C. Shephard. An elementary proof of Gram’s theorem for convex polytopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1214–1217, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Szasz:1967:PEG

[679] Paul Szász. On the pseudo-Euclidean geometry due to G. Hessenberg. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 1218–1232, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yaglom:1967:GMC

[680] I. M. Yaglom. Geometric models of certain genetic processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 1233–1242, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sion:1967:QM

[681] M. Sion and G. Zelmer. On quasimetrizability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1243–1249, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haimo:1967:EEH

[682] Franklin Haimo. Extensions of endomorphisms from the higher centres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):1250–1258, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Oppelt:1967:MAG

[683] John A. Oppelt. Mixed abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19 (??):1259–1262, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dyer:1967:TTC

[684] Micheal N. Dyer. Two term conditions in pi exact couples. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1263–1288, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Luh:1967:CR

[685] Jiang Luh. On the commutativity of J-rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1289–1292, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nelson:1967:PP

[686] R. Nelson. Pseudovaluations of polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1293–1302, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1967:SOD

[687] Homer G. Ellis. Solvability of ordinary differential equations near singular points: An analytic case. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 1303–1313, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Monroe:1967:GCE

[688] Itrel Monroe. On the generalized Cauchy equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1314–1318, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Watkins:1967:CCG

[689] M. E. Watkins and D. M. Mesner. Cycles and connectivity in graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??): 1319–1328, ???? 1967. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sonner:1967:LIP

[690] Johann Sonner. Lifting inductive and projective limits. Canadian Journal of Mathematics = Journal canadien de mathématiques, 19(??):1329–1339, ???? 1967. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaul:1968:ARH

[691] S. K. Kaul. On almost regular homeomorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1–6, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clunie:1968:IFH

[692] J. Clunie and T. Kövari. On integral functions having prescribed asymptotic growth. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):7–20, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlson:1968:CMQ

[693] R. E. Carlson and C. G. Cullen. Commutativity for matrices of quaternions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):21–24, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reilly:1968:BIS

[694] N. R. Reilly and A. H. Clifford. Bisimple inverse semigroups as semigroups of ordered triples. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):25–39, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Parthasarathy:1968:EGG

[695] K. R. Parthasarathy. Enumeration of graphs with given partition. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 40–47, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gard:1968:FDE

[696] J. R. Gard and R. D. Johnson. Four dimension equivalences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):48–50, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Agarwal:1968:PEF

[697] Arun Kumar Agarwal. On the properties of an entire function of two complex variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):51–57, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baker:1968:FVL

[698] Kirby A. Baker. Free vector lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):58–66, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Atkin:1968:RC

[699] A. O. L. Atkin. Ramanujan congruences for p_{−k}(n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):67–78, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Henstock:1968:RTI

[700] Ralph Henstock. A riemann-type integral of Lebesgue power. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 79–87, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clark:1968:GRR

[701] W. Edwin Clark. Generalized radical rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):88–94, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moon:1968:ICS

[702] J. W. Moon. On independent complete subgraphs in a graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):95–102, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dempster:1968:GCT

[703] M. A. H. Dempster. On the Gotlieb– Csima time-tabling algorithm. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 103–119, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Viswanathan:1968:GFU

[704] B. Viswanathan. Generating functions for ultra-spherical functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 120–134, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garg:1968:ADN

[705] K. M. Garg. On asymmetrical derivates of non-differentiable functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 135–143, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brualdi:1968:CSN

[706] R. A. Brualdi. Convex sets of nonnegative matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):144–157, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rigby:1968:AAG

[707] J. F. Rigby. Axioms for absolute geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):158–181, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Djabali:1968:AFA

[708] M. Djabali. Anneau de fractions d’un J-anneau. (French) [Ring of fractions of a J-ring]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):182–202, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thompson:1968:CMP

[709] R. C. Thompson. Commutators of matrices with prescribed determinant. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):203–221, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rajagopalan:1968:IMW

[710] M. Rajagopalan and K. G. Witz. On invariant means which are not inverse invariant. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):222–224, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Menon:1968:MLE

[711] M. V. Menon. Matrix links, an extremization problem, and the reduction of a non-negative matrix to one with prescribed row and column sums. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):

225–232, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Russo:1968:MSB

[712] John P. Russo. Monotone and eschauder bases of subspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 233–241, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Salomaa:1968:ERP

[713] Arto Salomaa. On events represented by probalistic automata of different types. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):242–251, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mahowald:1968:ASS

[714] Mark Mahowald. On the Adams spectral sequence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):252–256, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grace:1968:PCP

[715] E. E. Grace. Peripheral covering properties imply covering properties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 257–263, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tamura:1968:MGH

[716] Takayuki Tamura. Maximal or greatest homomorphic images of given type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):264–271, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rickart:1968:HCG

[717] C. E. Rickart. Holomorphic convexity for general function algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 272–290, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaufman:1968:PMA

[718] Robert Kaufman. Problems on measure algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):291–297, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kelly:1968:PZO

[719] John B. Kelly. Products of zeroone matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):298–329, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1968:VLD

[720] A. J. Ellis. Vector lattices with duals of integral type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):330–343, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bouwer:1968:SRS

[721] I. Z. Bouwer. Standard representations of simple Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):

344–361, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davison:1968:TTA

[722] T. M. K. Davison. A Tauberian theorem and analogues of the prime number theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):362–367, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robertson:1968:ODM

[723] James B. Robertson. Orthogonal decompositions of multivariate weakly stationary stochastic processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 368–383, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herstein:1968:SHG

[724] I. N. Herstein. Semi-homomorphisms of groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):384–388, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hager:1968:NCS

[725] Anthony W. Hager and Donald G. Johnson. A note on certain subalgebras of C(X). Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):389–393, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Srinivasacharyulu:1968:SNL

[726] K. Srinivasacharyulu. On some nonlinear problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):394–397, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muller:1968:CAD

[727] Bruno J. Müller. The classification of algebras by dominant dimension. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):398–409, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hill:1968:ABS

[728] J. D. Hill and W. T. Sledd. Approximation in bounded summability fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):410–415, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Patty:1968:HDPa

[729] C. W. Patty. Homology of deleted products of contractible 2-dimensional polyhedra. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):416–441, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nordgren:1968:CO

[730] Eric A. Nordgren. Composition operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):442–449, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Whitley:1968:SUS

[731] Robert Whitley. The size of the unit sphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):450–455, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Poland:1968:FGG

[732] John Poland. Finite groups with a given number of conjugate classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):456–464, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baxter:1968:RIP

[733] W. E. Baxter and W. S. Martindale III. Rings with involution and polynomial identities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):465–473, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Armendariz:1968:HPL

[734] E. P. Armendariz and W. G. Leavitt. The hereditary property in the lower radical construction. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):474–476, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Halperin:1968:NMI

[735] Israel Halperin. Von Neumann’s manuscript on inductive limits of regular rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):477–483, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorimer:1968:GGM

[736] P. J. Lorimer. On S_{2}-groups and groups of Moebius transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 484–485, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gross:1968:FEF

[737] Fred Gross. On factorization of elliptic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):486–494, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barnes:1968:GFT

[738] Bruce Alan Barnes. A generalized Fredholm theory for certain maps in the regular representations of an algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):495–504, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hungerford:1968:CHG

[739] Thomas W. Hungerford. A characterization of the hyperhomology groups of the tensor product. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):505–510, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Neumann:1968:CAS

[740] B. H. Neumann. Corrigendum and addendum to: “Some remarks on semigroup presentations”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):511, ????

1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [661].

Warne:1968:EEB

[741] R. J. Warne. Errata to: “Extensions of I-bisimple semigroups”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 511–512, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [602].

Dawkins:1968:CAI

[742] Brian P. Dawkins and Israel Halperin. Corrigendum and addendum to: “The isomorphism of certain continuous rings”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):512, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shapiro:1968:CMP

[743] H. L. Shapiro. Closed maps and paracompact spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):513–519, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sobczyk:1968:GCC

[744] Andrew Sobczyk. Graph-colouring and combinatorial numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 520–534, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dean:1968:NPO

[745] R. A. Dean and Gordon Keller. Natural partial orders. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):535–554, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wallach:1968:MSR

[746] Nolan R. Wallach. On maximal subsystems of root systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 555–574, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gratzer:1968:PAF

[747] George Grätzer. On polynomial algebras and free algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 575–581, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grillet:1968:FA

[748] Pierre-Antoine Grillet. Free P Qalgebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):582–595, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ewell:1968:DSS

[749] John A. Ewell. On the determination of sets by sets of sums of fixed order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):596–611, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leduc:1968:CSS

[750] P.-Y. Leduc. Catégories semi-simples et catégories primitives. (French)

[Semisimple categories and primitive categories]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):612–628, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Singh:1968:UEA

[751] K. D. Singh and N. D. Lane. Union and extension of arcs of cyclic order three. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):629–638, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rund:1968:GTM

[752] H. Rund. A geometrical theory of multiple integral problems in the calculus of variations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):639–657, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brooks:1968:ISL

[753] R. M. Brooks. On the invariance of the spectrum in locally m-convex algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):658–664, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kerber:1968:DK

[754] Adalbert Kerber. Zur Darstellungstheorie von Kranzprodukten. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 665–672, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lo:1968:CPS

[755] C.-Y. Lo. A class of polynomials in self-adjoint operators in spaces with an indefinite metric. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):673–678, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Greiner:1968:EGC

[756] Peter C. Greiner. Estimates for general coercive boundary problems on a halfspace for a class of elliptic partial differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):679–697, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hughes:1968:TCP

[757] J. W. B. Hughes. A theorem concerning partitions and its consequences in the theory of Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 698–700, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [1014].

Hoffman:1968:TVS

[758] Frederick Hoffman and Lloyd R. Welch. Totally variant sets in finite groups and vector spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):701–710, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Friesen:1968:FRG

[759] Donald K. Friesen. Finite rotation groups in low dimensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 711–719, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hallam:1968:SSO

[760] T. G. Hallam and V. Komkov. Stability of solutions of ordinary differential equations with respect to a closed set. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):720–726, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maddox:1968:MTI

[761] I. J. Maddox. Matrix transformations in an incomplete space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 727–734, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meir:1968:ITG

[762] A. Meir. An inclusion theorem for generalized Cesàro and Riesz means. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):735–738, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Botta:1968:LTM

[763] Peter Botta. Linear transformation on matrices: The invariance of a class of general matrix functions. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 739–748, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lynn:1968:DCS

[764] M. Stuart Lynn. On the diameter of a p-cyclic strongly connected digraph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):749–755, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nanzetta:1968:RTR

[765] Philip Nanzetta. A representation theorem for relatively complemented distributive lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):756–758, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoo:1968:MWI

[766] C. S. Hoo. Maps which induce the zero map on homotopy. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):759–768, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osborn:1968:CNA

[767] J. Marshall Osborn. Commutative non-associative algebras and identities of degree four. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):769–794, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borges:1968:MTS

[768] Carlos J. R. Borges. On metrizability of topological spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):795–804, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanRooij:1968:LAT

[769] A. C. M. van Rooij. The lattice of all topologies is complemented. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 805–807, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Read:1968:UFC

[770] Ronald C. Read. The use of Sfunctions in combinatorial analysis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):808–841, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Patty:1968:HDPb

[771] C. W. Patty. Homology of deleted products of contractible 2-dimensional polyhedra. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):842–854, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Knopp:1968:NCS

[772] Paul Knopp and Richard Sinkhorn. A note concerning simultaneous integral equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):855–861, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kruse:1968:RWA

[773] Robert L. Kruse. Rings in which all subrings are ideals. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):

862–871, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schmidt:1968:DSP

[774] Jürgen Schmidt. Direct sums of partial algebras and final algebraic structures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):872–887, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Guy:1968:CCG

[775] Richard K. Guy and Lowell W. Beineke. The coarseness of the complete graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):888–894, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osofsky:1968:ERQ

[776] B. L. Osofsky. Endomorphism rings of quasi-injective modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 895–903, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Derzko:1968:IRP

[777] N. A. Derzko. An integral representation for the product of spectral measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):904–912, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sterling:1968:RS

[778] Nicholas J. Sterling. Rings satisfying (x, y, z) = (y, z, x). Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):913–918, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gamelin:1968:EPS

[779] T. W. Gamelin and M. Voichick. Extreme points in spaces of analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):919–928, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Majindar:1968:SMC

[780] Kulendra N. Majindar. On some methods for construction of b. i. b. designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):929–938, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [956].

Trott:1968:ALG

[781] Stanton Trott. An axiomatic line geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):939–952, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Luther:1968:DM

[782] Norman Y. Luther. A decomposition of measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):953–959, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leviatan:1968:SMP

[783] D. Leviatan. Some moment problems in a finite interval. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):960–966, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Larman:1968:PUS

[784] D. G. Larman. On packings of unequal spheres in R_{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):967–969, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilmer:1968:CJO

[785] Robert Gilmer. On a condition of J. Ohm for integral domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 970–983, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Laudal:1968:PST

[786] Olav Arnfinn Laudal. Projective systems on trees and valuation theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):984–1000, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lowenthal:1968:SPG

[787] Franklin Lowenthal. On subsemigroups of the projective group on the line. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):1001–1011, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Radjavi:1968:SUI

[788] Heydar Radjavi. Simultaneous unitary invariants for sets of matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1012–1019, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Graham:1968:DM

[789] R. L. Graham and J. H. van Lint. On the distribution of nheta modulo 1. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1020–1024, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1968:STW

[790] Gillian M. Brown. A study of tensors which characterize a hypersurface of a Finsler space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1025–1036, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jones:1968:CCF

[791] William B. Jones and W. J. Thron. Convergence of continued fractions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):1037–1055, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1968:CRR

[792] Marvin Marcus and Stephen Pierce. On a combinatorial result of R. A. Brualdi and M. Newman. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1056–1067, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sheehan:1968:NGG

[793] J. Sheehan. The number of graphs with a given automorphism group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1068–1076, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Knuth:1968:AET

[794] Donald E. Knuth. Another enumeration of trees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1077–1086, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). Check issue number??

Thomas:1968:ARS

[795] J. Pelham Thomas. Associated regular spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1087–1092, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dixmier:1968:ELQ

[796] J. Dixmier. Sur les espaces localement quasi-compacts. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1093–1100, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glauberman:1968:CSS

[797] George Glauberman. A characteristic subgroup of a p-stable group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1101–1135, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rice:1968:MVL

[798] Norman M. Rice. Multiplication in vector lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1136–1149, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roy:1968:EPL

[799] Ashoke K. Roy. Extreme points and linear isometries of the Banach space of Lipschitz functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1150–1164, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gaudry:1968:IMA

[800] G. I. Gaudry. Isomorphisms of multiplier algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1165–1172, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodman:1968:VSP

[801] A. W. Goodman. The valence of sums and products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1173–1177, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pollak:1968:GLI

[802] Barth Pollak. Generation of local integral orthogonal groups in characteristic 2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):1178–1191, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1968:PFR

[803] Harvey A. Smith. Positive functionals and representations of tensor products of symmetric Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1192–1202, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nagabhushanam:1968:NSP

[804] K. Nagabhushanam and C. S. K. Bhagavan. Non-stationary processes and spectrum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1203–1206, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McWaters:1968:ASH

[805] M. M. McWaters. Arcs, semigroups, and hyperspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1207–1210, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zame:1968:ACW

[806] Alan Zame. Almost convergence and well-distributed sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1211–1214, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Magill:1968:RSC

[807] Kenneth D. Magill, Jr. Restrictive semigroups of closed functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):

1215–1229, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kelleher:1968:IMF

[808] James J. Kelleher. On isomorphisms of meromorphic function fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1230–1241, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barbeau:1968:SPS

[809] Edward J. Barbeau. Subordinate and pseudo-subordinate semi-algebras. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):1242–1245, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fox:1968:IEV

[810] Geoffrey Fox. Inductive extension of a vector measure under a convergence condition. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1246–1255, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hobby:1968:FGN

[811] C. Hobby. Finite groups with normal normalizers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1256–1260, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eakin:1968:SOQ

[812] Paul M. Eakin, Jr. and William J. Heinzer. Some open questions on minimal primes of a Krull domain. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1261–1264, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gross:1968:PCE

[813] Fred Gross. On the periodicity of compositions of entire functions. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1265–1268, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnsen:1968:SAG

[814] E. C. Johnsen. Special abelian group difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1269–1275, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gudder:1968:CSO

[815] Stanley P. Gudder. Complete sets of observables and pure states. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1276–1280, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodman:1968:SPO

[816] Ruth Goodman. Sets of polynomials orthogonal simultaneously on four ellipses. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1281–1294, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Priddis:1968:IBN

[817] M. J. M. Priddis. An inequality between numerical homotopy invariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):1295–1299, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gross:1968:SRG

[818] Fletcher Gross. Some remarks on groups admitting a fixed-point-free automorphism. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1300–1307, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Treybig:1968:CUS

[819] L. B. Treybig. Concerning upper semi-continuous decompositions of E^{n} whose non-degenerate elements are polyhedral arcs or star-like continua. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):1308–1314, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1968:FPC

[820] I. M. Isaacs. Fixed points and characters in groups with non-coprime operator groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1315–1320, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gupta:1968:ECG

[821] N. D. Gupta and M. F. Newman. Engel congruences in groups of primepower exponent. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1321–1323, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kwee:1968:STT

[822] B. Kwee. Some Tauberian theorems for the logarithmic method of summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1324–1331, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Majindar:1968:SPM

[823] Kulendra N. Majindar. On spectral properties of matrices with positive characteristic vectors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1332– 1343, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Klarner:1968:NCG

[824] David A. Klarner. The number of kcoloured graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1344–1352, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1968:OF

[825] Arlen Brown and Carl Pearcy. Operators of the form P AQ − QAP. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1353–1361, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Conrad:1968:LDS

[826] Paul F. Conrad. Lifting disjoint sets in vector lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1362–1364, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ton:1968:NLE

[827] Bui An Ton. Non-local elliptic boundary-value problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1365–1382, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Janos:1968:THC

[828] Ludvík Janos. Topological homotheties on compact metrizable spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1383–1386, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Janos:1968:LCH

[829] Ludvík Janos. Linearization of a contractive homeomorphism. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1387–1390, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramirez:1968:MAO

[830] Donald E. Ramirez. The measure algebra as an operator algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1391–1396, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alltop:1968:FPC

[831] W. O. Alltop. Free planes and collineations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1397–1411, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sallee:1968:PVE

[832] G. T. Sallee. Polytopes, valuations, and the Euler relation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1412–1424, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shapiro:1968:CSF

[833] Harold S. Shapiro. A class of singular functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1425–1431, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Behr:1968:PG

[834] H. Behr and J. Mennicke. A presentation of the groups PSL(2, p). Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1432–1438, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Block:1968:ELA

[835] Richard E. Block. On the extensions of Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1439–1450, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1968:GRF

[836] I. M. Isaacs and D. S. Passman. Groups with relatively few non-linear irreducible characters. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1451–1458, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kallaher:1968:RPP

[837] Michael J. Kallaher. Reidemeister projective planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1459–1464, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glauberman:1968:CCR

[838] George Glauberman. Correspondences of characters for relatively prime operator groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1465–1488, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pfeffer:1968:LDS

[839] W. F. Pfeffer. On the lower derivate of a set function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1489–1498, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Iltis:1968:SAS

[840] Richard Iltis. Some algebraic structure in the dual of a compact group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??): 1499–1510, ???? 1968. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miller:1968:CPG

[841] Donald J. Miller. The categorical product of graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1511–1521, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miller:1968:NIF

[842] John D. Miller. A note on involutions with a finite number of fixed points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20 (??):1522–1530, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoo:1968:HCL

[843] C. S. Hoo. On the homotopycommutativity of loop-spaces and suspensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):1531–1536, ???? 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Amitsur:1969:CAL

[844] S. A. Amitsur. Correction to “Arithmetic linear transformations”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):1–5, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [72].

Fraenkel:1969:BFC

[845] Aviezri S. Fraenkel. The bracket function and complementary sets of integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):6–27, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Puttaswamaiah:1969:URG

[846] B. M. Puttaswamaiah. Unitary representations of generalized symmetric groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):28–38, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DeMeyer:1969:PMC

[847] F. R. DeMeyer. Projective modules over central separable algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 39–43, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Childs:1969:PMA

[848] L. N. Childs. On projective modules and automorphismes of central separable algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):44–53, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Prasad:1969:APJ

[849] J. Prasad and A. K. Varma. An analogue of a problem of J. Balázs and P. Turán. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):54–63, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kantor:1969:CFP

[850] William M. Kantor. Characterizations of finite projective and affine spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):64–75, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Artmann:1969:HPD

[851] Benno Artmann. Hjelmslev planes derived from modular lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 76–83, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barnes:1969:FER

[852] Bruce Alan Barnes. The Fredholm elements of a ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):84–95, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deGroot:1969:CRS

[853] J. de Groot and J. M. Aarts. Complete regularity as a separation axiom. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):96–105, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lacroix:1969:TDL

[854] Norbert H. J. Lacroix. Twodimensional linear groups over local rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):106–135, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlson:1969:CIF

[855] R. E. Carlson and C. G. Cullen. A characterization of intrinsic functions on D_{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):136–146, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thomason:1969:NND

[856] S. K. Thomason. A note on nondistributive sublattices of degrees and hyperdegrees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):147–148, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Viswanathan:1969:GHT

[857] T. M. Viswanathan. Generalization of Hölder’s theorem to ordered modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):149–157, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Prugovecki:1969:TGI

[858] Eduard Prugovecki. Topologies on generalized inner product spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 158–169, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Antonelli:1969:STMa

[859] Peter L. Antonelli. Structure theory for Montgomery–Samelson fiberings between manifolds. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 170–179, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Antonelli:1969:STMb

[860] Peter L. Antonelli. Structure theory for Montgomery–Samelson fiberings between manifolds. I, II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 180–186, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Caveny:1969:ACF

[861] James Caveny. Absolute convergence factors for H^{p} series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):187–195, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kennison:1969:NCS

[862] J. F. Kennison. Normal completions of small categories. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):196–201, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lamoreaux:1969:DMS

[863] Jack W. Lamoreaux. Decomposition of metric spaces with a 0-dimensional set of non-degenerate elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 202–216, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sandler:1969:FCG

[864] Reuben Sandler. On finite collineation groups of F_{5}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):217–221, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beesack:1969:CDI

[865] Paul R. Beesack. On certain discrete inequalities involving partial sums. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):222–234, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lavie:1969:SDD

[866] Meira Lavie. The Schwarzian derivative and disconjugacy of nth order linear differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 235–249, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoo:1969:CS

[867] C. S. Hoo. On the criterion of Stasheff. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):250–255, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Atkin:1969:RC

[868] A. O. L. Atkin. Ramanujan congruences for p_{−k}(n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):256, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [898].

Sreenivasulu:1969:MTR

[869] V. Sreenivasulu. The maximum term and the rank of an entire function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):257–261, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [1013].

Rowlinson:1969:PCP

[870] Elizabeth Rowlinson and Hans Schwerdtfeger. Polynomials with certain prescribed conditions on the Galois group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):262–273, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1969:ESI

[871] W. G. Brown and J. W. Moon. Sur les ensembles de sommets indépendants dans les graphes chromatiques minimaux. (French) [On independent sets of vertices in minimum chromatic graphs]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):274–278, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wille:1969:VIM

[872] Rudolf Wille. Variety invariants for modular lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):279–283, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kelleher:1969:RMF

[873] James Kelleher. Rings of meromorphic functions on non-compact Riemann surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):284–300, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Michler:1969:IIP

[874] Gerhard Michler. Idempotent ideals in perfect rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):301–309, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zassenhaus:1969:PGT

[875] Hans J. Zassenhaus. A presentation of the groups PSL(2, p) with three defining relations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):310–311, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Caughran:1969:TRC

[876] James G. Caughran. Two results concerning the zeros of functions with finite Dirichlet integral. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):312– 316, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Milner:1969:PRO

[877] E. C. Milner. Partition relations for ordinal numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):317–334, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gorenstein:1969:FGC

[878] Daniel Gorenstein. Finite groups the centralizers of whose involutions have normal 2-complements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):

335–357, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Foulser:1969:CGG

[879] David A. Foulser. Collineation groups of generalized André planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 358–369, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brannan:1969:BST

[880] D. A. Brannan and Ch. Pommerenke. On the Bernstein–Szegő theorem for complex polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):370– 371, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anderson:1969:HRD

[881] T. Anderson. Hereditary radicals and derivations of algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 372–377, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cantor:1969:APT

[882] David G. Cantor. On arithmetic properties of the Taylor series of rational functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):378–382, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fugate:1969:CCC

[883] J. B. Fugate. A characterization of chainable continua. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):383–393, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Himmelberg:1969:SST

[884] C. J. Himmelberg and F. S. Van Vleck. Some selection theorems for measurable functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):394–399, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCandless:1969:OPS

[885] Byron H. McCandless. On order paracompact spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):400–405, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schafer:1969:AGV

[886] James A. Schafer. Abelian groups with a vanishing homology group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 406–409, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sehgal:1969:IIGa

[887] Sudarshan K. Sehgal. On the isomorphism of integral group rings. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 410–413, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Westwick:1969:LTG

[888] Roy Westwick. Linear transformations on Grassmann spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 414–417, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beidleman:1969:GFS

[889] James C. Beidleman. Generalized Frattini subgroups of finite groups. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):418–429, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Courter:1969:FDS

[890] Richard Courter. Finite direct sums of complete matrix rings over perfect completely primary rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 430–446, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gauthier:1969:CRA

[891] Paul Gauthier. Cercles de remplissage and asymptotic behaviour. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 447–455, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gray:1969:IIA

[892] Alfred Gray. Isometric immersions of almost Hermitian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):

456–459, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kreith:1969:LBE

[893] Kurt Kreith. Lower bounds for the essential spectrum of fourth-order differential operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):460–465, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Watters:1969:LRA

[894] J. F. Watters. Lower radicals in associative rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):466–476, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Werner:1969:DCT

[895] W. L. Werner. A double-centralizer theorem for simple associative algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):477–478, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoo:1969:HNL

[896] C. S. Hoo. Homotopical nilpotency of loop-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):479–484, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1969:ISM

[897] Marvin Marcus and Paul J. Nikolai. Inequalities for some monotone matrix functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):485–494, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beasley:1969:RCC

[898] LeRoy B. Beasley. Ramanujan congruences for p_{−k}(n): Corrigendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 495–497, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [868].

Derderian:1969:GCP

[899] J.-C. Derdérian. Galois connections and pair algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):498–501, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gray:1969:FPT

[900] William J. Gray. A fixed-point theorem for commuting monotone functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):502–504, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stampfli:1969:AAO

[901] J. G. Stampfli. Adjoint abelian operators on Banach space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 505–512, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zamfirescu:1969:PCF

[902] Tudor Zamfirescu. On planar continuous families of curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 513–530, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kerr-Lawson:1969:SLI

[903] A. Kerr-Lawson. Some lemmas on interpolating Blaschke products and a correction. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):531–534, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mohanty:1969:BSA

[904] R. Mohanty and B. K. Ray. On the behaviour of a series associated with the conjugate series of a Fourier series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):535–551, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mohanty:1969:ACS

[905] R. Mohanty and B. K. Ray. On the absolute Cesàro summability of negative order of a series associated with the conjugate series of a Fourier series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):552–557, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arnold:1969:ITK

[906] Jimmy T. Arnold. On the ideal theory of the Kronecker function ring and the domain D(X). Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):558–563, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Christiansen:1969:RTG

[907] R. A. Christiansen. Recurrent transformation groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):564–575, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ohm:1969:SVG

[908] Jack Ohm. Semi-valuations and groups of divisibility. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):576–591, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ruston:1969:NCC

[909] A. F. Ruston. A note on the Caradus class F of bounded linear operators on a complex Banach space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 592–594, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

MacLane:1969:GBP

[910] G. R. MacLane and L. A. Rubel. On the growth of Blaschke products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 595–601, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garling:1969:CRS

[911] D. J. H. Garling. A class of reflexive symmetric BK-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 602–608, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Seo:1969:GGC

[912] Tae Kun Seo and G. Whaples. A generalization of global class field theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):609–614, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wenger:1969:SSH

[913] R. Wenger. Some semigroups having quasi-Frobenius algebras. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 615–624, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Keown:1969:BSS

[914] R. Keown and C. Conatser. Biorthogonal systems in l^{p}-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 625–638, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lardy:1969:MRI

[915] L. J. Lardy and J. A. Lindberg, Jr. On maximal regular ideals and identities in the tensor product of commutative Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):639–647, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:1969:SLO

[916] Peng yee Lee. Some linear operators in the L^{p} spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):648–654, ????

1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramsay:1969:GEU

[917] R. T. Ramsay. Groups with equal uniformities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):655–659, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Timourian:1969:MDB

[918] J. G. Timourian. Maps with discrete branch sets between manifolds of codimension one. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):660–668, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hahn:1969:GSP

[919] K. T. Hahn and Josephine Mitchell. Generalization of Schwarz–Pick lemma to invariant volume. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):669–674, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stolarsky:1969:LBS

[920] Kenneth B. Stolarsky. A lower bound for the Scholz–Brauer problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 675–683, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1969:MST

[921] Benson Samuel Brown. The mod C suspension theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):684–701, ????

1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1969:FA

[922] Benson Samuel Brown. A first approximation to X, Y. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):702–711, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1969:SSC

[923] Benson Samuel Brown. A spectral sequence for cohomotopy. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 712–729, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoo:1969:CWC

[924] C. S. Hoo. Conilpotency and weak category. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):730–734, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tymchatyn:1969:OPS

[925] E. D. Tymchatyn. A one-parameter subsemigroup which meets many regular D-classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):735–739, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1969:TTB

[926] D. Borwein. A Tauberian theorem for borel-type methods of summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):740–747, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wenner:1969:RSC

[927] B. R. Wenner. Remetrization in strongly countable-dimensional spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):748–750, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mullins:1969:CAF

[928] Robert E. Mullins. A characterization of the algebra of functions vanishing at infinity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):751–754, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rayburn:1969:LA

[929] Marlon C. Rayburn. On the lattice of σ-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):755–761, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hill:1969:DG

[930] Paul Hill. On the decomposition of groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):762–768, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bender:1969:ESC

[931] Edward A. Bender and Thomas W. Tucker. The exponents of strongly connected graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):769–782, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohn:1969:SI

[932] Richard M. Cohn. Systems of ideals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):783–807, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kinney:1969:LRP

[933] J. R. Kinney and T. S. Pitcher. On the lower range of Perron’s modular function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):808–816, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lozier:1969:CCR

[934] F. W. Lozier. A class of compact rigid 0-dimensional spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 817–821, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elliott:1969:DPR

[935] P. D. T. A. Elliott. The distribution of primitive roots. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):822–841, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dyer:1969:FCC

[936] Micheal Dyer. Federer–Čech couples. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):842–864, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burgess:1969:RQG

[937] W. D. Burgess. Rings of quotients of group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):865–875, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rigby:1969:AAG

[938] J. F. Rigby. Axioms for absolute geometry. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):876–883, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:1969:SLC

[939] C. C. Chen and G. Grätzer. Stone lattices. I. construction theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 884–894, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:1969:SLI

[940] C. C. Chen and G. Grätzer. Stone lattices. II. structure theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 895–903, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lanski:1969:NSG

[941] Charles Lanski. Nil subrings of Goldie rings are nilpotent. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):904–907, ????

1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Humphreys:1969:AIC

[942] J. E. Humphreys. On the automorphisms of infinite Chevalley groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):908–911, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wilken:1969:NSR

[943] Donald R. Wilken. A note on strongly regular function algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 912–914, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gewirtz:1969:GME

[944] A. Gewirtz. Graphs with maximal even girth. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):915–934, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaul:1969:TG

[945] S. K. Kaul. On a transformation group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):935–941, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nasim:1969:SRR

[946] C. Nasim. Some results relating the behaviour of Fourier transforms near the origin and at infinity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):

942–950, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nasim:1969:SFI

[947] C. Nasim. A summation formula involving σ_{k}(n), k > 1. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):951–964, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Billis:1969:FGW

[948] M. J. Billis. Finite groups whose powers have no countably infinite factor groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):965–969, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roggenkamp:1969:ILO

[949] Klaus W. Roggenkamp. On the irreducible lattices of orders. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 970–976, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Labelle:1969:RMA

[950] G. Labelle and Q. I. Rahman. Remarque sur la moyenne arithmétique de fonctions univalentes convexes. (French) [Note on the arithmetic mean of convex univalent functions]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 977–981, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nikolai:1969:MVL

[951] Paul J. Nikolai. Mean value and limit theorems for generalized matrix functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):982–991, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beineke:1969:MDC

[952] L. W. Beineke. Minimal decompositions of complete graphs into subgraphs with embeddability properties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 992–1000, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eaton:1969:NAL

[953] W. T. Eaton. A note about locally spherical spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1001–1003, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Byrd:1969:ACL

[954] Richard D. Byrd. Archimedean closures in lattice-ordered groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 1004–1012, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wyman:1969:AB

[955] M. Wyman and R. Wong. The asymptotic behaviour of μ(z, β, α). Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):

1013–1023, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Majindar:1969:SMC

[956] Kulendra N. Majindar. “On some methods of construction of b. i. b. designs”: Corrigendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1024, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [780].

Wales:1969:FLGa

[957] David B. Wales. Finite linear groups of prime degree. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1025–1041, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wales:1969:FLGb

[958] David B. Wales. Finite linear groups of degree seven. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1042–1056, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1969:PIF

[959] William W. Smith. Projective ideals of finite type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1057–1061, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bercov:1969:STP

[960] R. D. Bercov. On simply transitive primitive permutation groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):

1062–1068, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sauer:1969:EPE

[961] N. Sauer. Extremal point and edge sets in n-graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1069–1075, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dubuc:1969:TAA

[962] Serge Dubuc. Transformations affines agissant sur un ensemble compact convexe. (French) [Affine transformations acting on a compact convex set]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 1076–1082, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kirkpatrick:1969:NCP

[963] P. B. Kirkpatrick. A note on the construction of projective planes from groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1083–1085, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beineke:1969:CCB

[964] Lowell W. Beineke and Richard K. Guy. The coarseness of the complete bipartite graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1086–1096, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benz:1969:GKG

[965] Walter Benz. Zur Geometrie der Körpererweiterungen. (German) [The geometry of field extensions]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 1097–1122, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rhodes:1969:HGT

[966] F. Rhodes. Homotopy groups of transformation groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1123–1136, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Passman:1969:SMG

[967] D. S. Passman. On the semisimplicity of modular group algebras. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 1137–1145, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putz:1969:CTF

[968] H. Putz. Construction of transverse fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1146–1159, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rhemtulla:1969:CCF

[969] A. H. Rhemtulla. Commutators of certain finitely generated soluble groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):1160–1164, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wenzel:1969:AGU

[970] G. H. Wenzel. Automorphism groups of unary algebras on groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 1165–1171, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

MacGregor:1969:ICA

[971] Thomas H. MacGregor. An inequality concerning analytic functions with a positive real part. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1172–1177, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davis:1969:OAI

[972] Chandler Davis, Heydar Radjavi, and Peter Rosenthal. On operator algebras and invariant subspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 1178–1181, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sehgal:1969:IIGb

[973] Sudarshan K. Sehgal. On the isomorphism of integral group rings. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 1182–1188, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Finkelstein:1969:RPB

[974] Mark Finkelstein. The reflection principle for Banach space-valued analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1189–1191, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Megibben:1969:GKC

[975] Charles Megibben. The generalized Kulikov criterion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1192–1205, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McArthur:1969:CAB

[976] C. W. McArthur, Ivan Singer, and Mark Levin. On the cones associated with biorthogonal systems and bases in Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1206–1217, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaplansky:1969:AQF

[977] Irving Kaplansky and Richard J. Shaker. Abstract quadratic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):1218–1233, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Camina:1969:GMS

[978] A. R. Camina and T. M. Gagen. Groups with metacyclic Sylow 2subgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1234–1237, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alperin:1969:CAG

[979] J. L. Alperin. A classification of nabelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1238–1244, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boyd:1969:IFS

[980] David W. Boyd. Indices of function spaces and their relationship to interpolation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1245–1254, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barth:1969:QSC

[981] K. F. Barth and W. J. Schneider. On a question of Seidel concerning holomorphic functions bounded on a spiral. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1255–1262, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edge:1969:TPS

[982] W. L. Edge. Three plane sextics and their automorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1263– 1278, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Howard:1969:OCM

[983] Henry C. Howard. “oscillation criteria for matrix differential equations”: Corrigendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1279–1280, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [579].

Shauck:1969:AHF

[984] Maxwell E. Shauck. Algebras of holomorphic functions in ringed spaces. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):1281–1292, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ching:1969:NIN

[985] Wai mee Ching. Non-isomorphic nonhyperfinite factors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1293–1308, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stewart:1969:UPD

[986] James Stewart. Unbounded positive definite functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1309–1318, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blatt:1969:SHM

[987] D. Blatt and G. Szekeres. A skew Hadamard matrix of order 52. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 1319–1322, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Welsh:1969:TTM

[988] D. J. A. Welsh. Transversal theory and matroids. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1323–1330, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Groemer:1969:PSP

[989] H. Groemer. On plane sections and projections of convex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 1331–1337, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muller:1969:MD

[990] Bruno J. Müller. On Morita duality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):1338–1347, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chacron:1969:TH

[991] M. Chacron. On a theorem of Herstein. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):1348–1353, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Geller:1969:RC

[992] Dennis Geller and Bennet Manvel. Reconstruction of cacti. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1354– 1360, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koch:1969:SRC

[993] C. F. Koch. Some remarks concerning (f, d_{n}) and [F, d_{n}] summability methods. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1361–1365, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ligh:1969:DNR

[994] Steve Ligh. On division near-rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):1366–1371, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1969:CST

[995] Norman Lloyd Johnson. A classification of semi-translation planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 1372–1387, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lyndon:1969:GGT

[996] R. C. Lyndon and J. L. Ullman. Groups generated by two parabolic linear fractional transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1388– 1403, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brungs:1969:GDV

[997] H.-H. Brungs. Generalized discrete valuation rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1404–1408, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kallaher:1969:RBQ

[998] Michael J. Kallaher. Right bol quasi-fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1409–1420, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Radjavi:1969:SAF

[999] Heydar Radjavi. On self-adjoint factorization of operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1421–1426, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McMullen:1969:LSP

[1000] P. McMullen. Linearly stable polytopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1427–1431, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moody:1969:ELA

[1001] Robert V. Moody. Euclidean Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1432–1454, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stephenson:1969:UAR

[1002] W. Stephenson. Unique addition rings.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):1455–1461, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rigby:1969:POP

[1003] J. F. Rigby. Pascal ovals in projective planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1462–1476, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robertson:1969:CFB

[1004] M. S. Robertson. Coefficients of functions with bounded boundary rotation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):1477–1482, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Williams:1969:DVP

[1005] Kenneth S. Williams. Distinct values of a polynomial in subsets of a finite field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):1483–1488, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Antonelli:1969:DMS

[1006] Peter L. Antonelli. Differentiable Montgomery–Samelson fiberings with finite singular sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1489–1495, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Douglas:1969:GCO

[1007] A. J. Douglas. Generalized Casimir operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1496–1505, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hedrlin:1969:CGG

[1008] Z. Hedrlín and E. Mendelsohn. The category of graphs with a given subgraph—with applications to topology and algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??):1506–1517, ????

1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Holland:1969:PSM

[1009] Samuel S. Holland, Jr. Partial solution to Mackey’s problem about modular pairs and completeness. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 1518–1525, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haddix:1969:PIE

[1010] G. F. Haddix, J. N. Mordeson, and B. Vinograde. On purely inseparable extensions of unbounded exponent. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21 (??):1526–1532, ???? 1969. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Choe:1969:NLC

[1011] Tae Ho Choe. Notes on locally compact connected topological lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 21(??): 1533–1536, ???? 1969. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shrikhande:1970:NES

[1012] S. S. Shrikhande, D. Raghavarao, and S. K. Tharthare. “Non-existence of some unsymmetrical partially balanced incomplete block designs”: Corrigendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 22(??):191, ???? 1970. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See

Sreenivasulu:1971:MTR

[1013] V. Sreenivasulu. “The maximum term and the rank of an entire function”: Corrigendum. Canadian Journal of Mathematics = Journal canadien de mathématiques, 23(??):131, ???? 1971. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [869].

Hughes:1973:CTC

[1014] J. W. B. Hughes. Corrigendum: “A theorem concerning partitions and its consequence in the theory of Lie algebras”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 25(??):224, ???? 1973. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [757].